### v2

# Real projective groups are formal

Abstract.
We prove that the mod $2$ cohomology algebras of real projective groups are intrinsically formal. As a consequence we derive the Hopkins–Wickelgren formality and the strong Massey vanishing conjecture for a large class of fields.
1991 Mathematics Subject Classification:
12G05, 16E40, 18N40, 55S30.
1. Introduction
Let $G$ be a profinite group. An embedding problem for $G$ is a solid diagram:
[…]
where $H$ and $\widetilde{H}$ are finite groups, the solid arrows are continuous homomorphisms and $\alpha$ is surjective. A solution of this embedding problem is a continuous homomorphism $\widetilde{\phi}\colon G\to\widetilde{H}$ which makes the diagram commutative. We say that the embedding problem above is real if for every involution $t\in G$ with $\phi(t)\neq 1$ there is an involution $\tilde{h}\in\widetilde{H}$ with $\alpha(\tilde{h})=\phi(t)$, i.e., if involutions do not provide an obstruction for the existence of a solution.
Definition 1.1.
Following Haran and Jarden [11] we say that a profinite group $G$ is real projective if $G$ has an open subgroup without $2$-torsion, and if every real embedding problem for $G$ has a solution.
For a prime number $p$ and a profinite group $G$, let $C^{*}(G,\mathbb{F}_{p})$ denote the differential graded $\mathbb{F}_{p}$-algebra of continuous cochains of $G$ with values in $\mathbb{F}_{p}$. The aim of this paper is to show that the differential graded algebra $C^{*}(G,\mathbb{F}_{p})$ of a real projective group $G$ has the following strong property.
Definition 1.2.
Let $\mathbb{F}$ be a field and let $C^{*}$ be a differential graded $\mathbb{F}$-algebra with cohomology $H^{*}$. Then $C^{*}$ is called formal if there is a sequence $C^{*}\leftarrow T^{*}\rightarrow H^{*}$ of quasi-isomorphisms of differential graded algebras between $C^{*}$ and $H^{*}$ where we consider $H^{*}$ as a differential graded $\mathbb{F}$-algebra with trivial differential. Moreover, a graded $\mathbb{F}$-algebra $A$ is called intrinsically formal if every differential graded algebra $C^{*}$ with $H^{*}(C^{*})\cong A$ is formal.
The main result of this paper is the following:
Theorem 1.3.
Let $G$ be a real projective profinite group. Then the graded algebra $H^{*}(G,\mathbb{F}_{2})$ is intrinsically formal. In particular, the differential graded $\mathbb{F}_{2}$-algebra $C^{*}(G,\mathbb{F}_{2})$ is formal.
Remark 1.4.
For an odd prime $p$, the $\mathbb{F}_{p}$-algebra $H^{*}(G,\mathbb{F}_{p})$ is concentrated in degrees 00 and $1$, i.e., $H^{i}(G,\mathbb{F}_{p})=0$ for $i\geq 2$. The intrinsic formality of $H^{*}(G,\mathbb{F}_{p})$ then follows easily from Corollary 4.9 and the argument of the proof of Lemma 6.5 below. Hence, together with Theorem 1.3, we can conclude that $H^{*}(G,\mathbb{F}_{p})$ is intrinsically formal and therefore $C^{*}(G,\mathbb{F}_{p})$ is formal for all primes $p$.
Our proof of Theorem 1.3 relies on a theorem due to Kadeishvili which states that a positively graded algebra $A$ over a field is intrinsically formal if the Hochschild cohomology groups $\mathrm{HH}^{n,2-n}(A,A)$ vanish for all $n\geq 3$. In order to show this vanishing for $A=H^{*}(G,\mathbb{F}_{2})$ we deduce from Scheiderer’s work in [33] the description of $H^{*}(G,\mathbb{F}_{2})$ for a real projective group $G$ as the following type of quadratic algebra: First, let $B$ be a Boolean ring, i.e., $x^{2}=x$ for every $x\in B$, and let $B_{*}=\bigoplus_{j\geq 0}B_{j}$ be its associated Boolean graded algebra, i.e., the $\mathbb{F}_{2}$-algebra with $B_{0}=\mathbb{F}_{2}$ and, for every positive integer $j$, $B_{j}=B$. Second, for an $\mathbb{F}_{2}$-vector space $V$, let $V_{*}=\bigoplus_{j\geq 0}V_{j}$ denote the graded algebra with $V_{0}=\mathbb{F}_{2}$, $V_{1}=V$ and $V_{j}=0$ for $j\geq 2$. We refer to Section 6 and Definitions 6.1 and 6.4 for more details. We show that $H^{*}(G,\mathbb{F}_{2})$ is a quadratic algebra given as the connected sum of quadratic algebras $B_{*}\sqcap V_{*}$ (see Definition 5.2) where $V$ is a suitable vector space and $B$ is the Boolean ring of continuous functions from the space of conjugacy classes of involutions in $G$ to $\mathbb{F}_{2}$. Then we calculate the Hochschild cohomology groups of such an algebra by an explicit description of the cocycles and coboundaries in the case of a locally finite Boolean graded algebra in Theorem 7.13. Finally, we use a spectral sequence argument and the vanishing of the higher derived limits of certain non-vanishing Hochschild cohomology groups to deduce the general case. Our main technical result is the following, see Theorem 9.14:
Theorem 1.5.
Let $B$ be a Boolean ring with $|B|\geq 8$, and let $V$ be an $\mathbb{F}_{2}$-vector space. Then
$\mathrm{HH}^{k,m}(B_{*}\sqcap V_{*},B_{*}\sqcap V_{*})=0$
for every $m<0$ and $k\geq 0$ such that $k\neq 1-m$.
Remark 1.6.
In fact, we thereby prove that for every profinite group $G$ such that $H^{*}(G,\mathbb{F}_{2})$ is of the form $B_{*}\sqcap V_{*}$ as in Theorem 1.5, the graded algebra $H^{*}(G,\mathbb{F}_{2})$ is intrinsically formal. However, real projective groups are the main examples of groups with this type of cohomology algebra that we are aware of.
One of the main consequences of formality of a differential graded associative algebra is the vanishing of Massey products. We will formulate a stronger form of the latter next.
Definition 1.7.
Let $\mathbb{F}$ be a field and let $(C^{*},\cup,\delta)$ be a differential graded associative $\mathbb{F}$-algebra with differential $\delta$. For an element $a\in C^{d}$ we write $\bar{a}:=(-1)^{1+d}a$. For an integer $n\geq 2$, let $a_{1},a_{2},\ldots,a_{n}$ be a set of cohomology classes with $a_{i}\in H^{d_{i}}$. A defining system for the $n$-fold Massey product of $a_{1},a_{2},\ldots,a_{n}$ is a set $\{a_{i,j}\}$ of elements of $C^{d_{i,j}}$ for $11\leq i<j\leq n+$ 1 and $(i,j)\neq(1,n+1)$ such that
$\delta(a_{i,j})=\sum_{k=i+1}^{j-1}\bar{a}_{i,k}\cup a_{k,j}$
and $a_{1},a_{2},\ldots,a_{n}$ is represented by $a_{1,2},a_{2,3},\ldots,a_{n,n+1}$. The degree $d_{i,j}$ of $a_{i,j}$ satisfies
$d_{i,j}=\sum_{s=i}^{j-1}d_{s}-(j-1-i)~{}\text{for}~{}i+1\leq j\leq n+1.$
We say that the $n$-fold Massey product of $a_{1},a_{2},\ldots,a_{n}$ is defined if there exists a defining system. The $n$-fold Massey product $\langle a_{1},a_{2},\ldots,a_{n}\rangle_{\{a_{i,j}\}}$ of $a_{1},a_{2},\ldots,a_{n}$ with respect to the defining system $\{a_{i,j}\}$ is the cohomology class of
$\sum_{k=2}^{n}\bar{a}_{1,k}\cup a_{k,n+1}$
in $H^{d_{n+1}}$ where $d_{n+1}=\sum_{i=1}^{n}d_{i}-n+2$. Let $\langle a_{1},a_{2},\ldots,a_{n}\rangle$ denote the subset of $H^{d_{n+1}}$ consisting of the $n$-fold Massey products of $a_{1},a_{2},\ldots,a_{n}$ with respect to all defining systems. We say that the $n$-fold Massey product of $a_{1},a_{2},\ldots,a_{n}$ vanishes if $\langle a_{1},a_{2},\ldots,a_{n}\rangle$ contains zero.
We say that $C^{*}$ satisfies strong Massey vanishing if for very $a_{1},a_{2},\ldots,a_{n}$ as above the following assertion holds: if
(1) $a_{1}\cup a_{2}=a_{2}\cup a_{3}=\cdots=a_{n-1}\cup a_{n}=0,$
then the $n$-fold Massey product of $a_{1},a_{2},\ldots,a_{n}$ vanishes. When condition (1) above is satisfied we say that all neighbouring cup products vanish for the $n$-tuple $a_{1},a_{2},\ldots,a_{n}$.
For real projective groups, strong Massey vanishing for cohomology classes in degree one was proven in [26, Theorem 1.7]. For pro $p$-groups of elementary type, strong Massey vanishing for cohomology classes in degree one was proven in [32, Theorem 1.2]. (See also the comments below Theorem 1.10.) As we explain in section 3, Theorem 1.3 implies the following result:
Theorem 1.8.
Let $G$ be a real projective profinite group, and let $p$ be a prime number. Then $C^{*}(G,\mathbb{F}_{p})$ satisfies strong Massey vanishing for cohomology classes in arbitrary degrees.
Remark 1.9.
For an odd prime $p$, $C^{*}(G,\mathbb{F}_{p})$ satisfies strong Massey vanishing since $H^{i}(G,\mathbb{F}_{p})=0$ for $i\geq 2$. Hence the difficult case is $p=2$. We emphasise that Theorem 1.8 proves the strong Massey vanishing conjecture for real projective groups for cohomology classes in arbitrary degrees. We thereby provide the first nontrivial class of profinite groups for which strong Massey vanishing holds beyond degree one.
Our interest in the formality of profinite groups originates from the arithmetic of fields. Let $F$ be any field whose characteristic is not two, and let $\Gamma(F)$ denote its absolute Galois group. In [13, Question 1.4 on page 1306] Hopkins and Wickelgren raised the question whether $C^{*}(\Gamma(F),\mathbb{F}_{2})$ is formal. We know that the answer is no in general. In fact, Positselski proved in [28, Section 9.11] that, for an odd prime number $p$ and a finite extension $F$ of $\mathbb{Q}_{p}$ which contains a primitive $p$-root of unity, the differential graded algebra $C^{*}(\Gamma(F),\mathbb{F}_{2})$ is not formal (see also [30, §6]). We note that there is a counterexample to the strong Massey vanishing for fourfold products over $\mathbb{Q}$, due to Wittenberg–Harpaz, in [9, Example A.15]. This implies that the absolute Galois group of $\mathbb{Q}$ is not formal. In the more recent paper [19] Merkurjev–Scavia provide a counterexample over a field containing the algebraic closure of $\mathbb{Q}$, so even a more restricted prediction of Positselski on the formality of fields is false, see [19, Question 1.5 and Theorem 1.6] and [20]. However, we still expect that many fields have formal absolute Galois groups. Our result is actually the first affirmative contribution to this conjecture as we will explain next.
Recall that a field $F$ has virtual cohomological dimension $\leq 1$ if there is a finite extension $K/F$ with $\mathrm{cd}(\Gamma(K))\leq 1$. As we recall in section 2, it is a consequence of classical Artin–Scheier theory and Haran’s work that the absolute Galois group of a field with virtual cohomological dimension $\leq 1$ is real projective. Hence our results for real projective groups imply the following:
Theorem 1.10.
Let $F$ be a field with virtual cohomological dimension $\leq 1$, and let $p$ be a prime number. Then $H^{*}(F,\mathbb{F}_{p})$ is intrinsically formal, and hence $C^{*}(\Gamma(F),\mathbb{F}_{p})$ is formal and satisfies strong Massey vanishing.
This theorem is also a new contribution to the Massey vanishing conjecture by Mináč and Tân [21, Conjecture 1.1], a very active field of research (see e.g. [7], [9], [13], [19], [22], [26], [32]). In fact, our theorem shows Massey vanishing in arbitrary cohomology degrees for fields with virtual cohomological dimension $\leq 1$. This is a much stronger result than in the previously known cases where only the degree one Massey products were considered.
Another consequence of our methods is a positive case of a conjecture by Positselski and Voevodsky on the Koszulity of Galois cohomology [29, §0.1, page 128]. In [24], other positive cases of this conjecture were proven. Recall that a positively graded algebra $A$ over $\mathbb{F}_{p}$ is called Koszul if the groups $H_{i,j}(A)=\mathrm{Tor}^{A}_{i,j}(\mathbb{F}_{p},\mathbb{F}_{p})$ vanish for all $i\neq j$, where the first grading $i$ is the homological grading and the second grading $j$ is the internal grading which is induced from the grading of $A$. Based on Scheiderer’s structure theorem for real projective groups we prove:
Theorem 1.11.
Let $F$ be a field with virtual cohomological dimension $\leq 1$, and let $p$ be a prime number. Then the cohomology algebra $H^{*}(F,\mathbb{F}_{p})$ is Koszul.
Contents.
In Section 2 we review the work of Haran–Jarden who showed that every real projective group arises as an absolute Galois group of a pseudo real closed field and deduce the structure of the mod $2$ cohomology algebra of real projective groups from Scheiderer’s local-global principle for real projective groups. In Section 3 we show that formality implies strong Massey vanishing for differential graded algebras. In Section 4 we review a Kadeishvili’s sufficient condition for intrinsic formality in terms of graded Hochschild cohomology. In Section 5 we recall the Koszul complex of quadratic algebras and the notion of Koszul algebras. In Section 6 we introduce Boolean and dual graded algebras and show that their connected sum is Koszul. In Section 7 we prove our main technical result on the graded Hochschild cohomology of the connected sum of dual and Boolean graded algebras, first for finite Boolean subrings. In Section 8 we recall some facts needed on Mittag-Leffler functors. In Section 9 we prove the main technical result Theorem 1.5. then the infinite Boolean case in Section 9. In Section 10 we combine the results of the previous sections to prove Theorem 1.3 and its consequences. In Appendix A we provide proofs of the assertions on Boolean rings we use in Sections 6, 7 and 9. In Appendix B we prove the existence of the spectral sequence that we use in section 9.
Acknowledgement.
The first-named author wishes to thank the hospitality of the Departments of Mathematical Sciences at University of Oslo and NTNU in Trondheim. Both authors gratefully acknowledge the partial support by the RCN Frontier Research Group Project No. 250399 Motivic Hopf equations and the RCN Project No. 313472 Equations in Motivic Homotopy. We thank the anonymous referee for many comments and suggestions which helped improving the article.
2. Fields whose absolute Galois group is real projective and Scheiderer’s theorem on their Galois cohomology
In this section, we first review the work of Haran–Jarden who showed that every real projective group arises as an absolute Galois group of a pseudo real closed field. Then we deduce from Scheiderer’s local-global principle the structure of the mod $2$ cohomology algebra of real projective groups.
For a profinite group $G$, let $\mathrm{cd}(G)$ denote its cohomological dimension as defined in [36]. Haran gave the following very useful characterisation of real projective groups [10, Proposition 2.2 on page 226]:
Theorem 2.1 (Haran).
A profinite group $G$ is real projective if and only if $G$ has an open subgroup $G_{0}$ of index $\leq 2$ with $\mathrm{cd}(G_{0})\leq 1$, and every involution $t\in G$ is self-centralising, that is, we have $C_{G}(t)=\{1,t\}$.
Haran’s theorem immediately implies that every closed subgroup of a real projective group $G$ is also real projective, and in particular every torsion element of $G$ has order dividing $2$.
For a field $F$, let $\Gamma(F)$ denote its absolute Galois group, and we write $\mathrm{cd}(F)=\mathrm{cd}(\Gamma(F)$. Recall that a field $F$ has virtual cohomological dimension $\leq 1$ if there is a finite extension $K/F$ with $\mathrm{cd}(K)\leq 1$. Since the only torsion elements in the absolute Galois group of $F$ are the involutions coming from the orderings of $F$, it is equivalent (by a theorem in [37]) to require $\mathrm{cd}(K)\leq 1$ for any fixed finite separable extension $K$ of $F$ without orderings, for example for $K=F(\mathbf{i})$, where $\mathbf{i}=\sqrt{-1}$. In particular, if $F$ itself cannot be ordered (which is equivalent to $-1$ being a sum of squares in $F$), this condition is equivalent to $\mathrm{cd}(F)\leq 1$.
Examples 2.2.
Examples of fields $F$ which can be ordered with $\mathrm{cd}(F(\mathbf{i}))\leq 1$ include real closed fields, function fields in one variable over any real closed ground field, the field of Laurent series in one variable over any real closed ground field, and the field $\mathbb{Q}^{ab}\cap\mathbb{R}$ which is the subfield of $\mathbb{R}$ generated by the numbers $\cos(\frac{2\pi}{n})$, where $n\in\mathbb{N}$.
The following is a well-known consequence of Artin–Schreier theory.
Lemma 2.3.
Let $\Gamma$ be an absolute Galois group and let $x\in\Gamma$ be an involution. Then the centraliser $C_{\Gamma}(x)$ is the group $C=\{1,x\}$ generated by $x$.
Proof.
Let $K$ be a field whose absolute Galois group is isomorphic to $\Gamma$ and let $\overline{K}$ be the separable closure of $K$. Let $R\subset\overline{K}$ be the fixed field of $C$. Then the absolute Galois group of $R$ is $\mathrm{Gal}(\overline{K}/R)=C\equiv\mathbb{Z}/2$, so $R$ is real closed by [18, Proposition 2.4, page 452]. In particular, $R$ has characteristic zero, so the same holds for all subfields of $\overline{K}$. Moreover, the latter is algebraically closed. Since $C_{\Gamma}(x)$ is the pre-image of $1$ under the map $y\mapsto x^{-1}y^{-1}xy$, it is closed. Let $F\subset\overline{K}$ be the fixed field of $C_{\Gamma}(x)$. Since $C\subseteq C_{\Gamma}(x)$ we get that $F$ is a subfield of $R$. Note that $C_{\Gamma}(x)$ leaves $R$ invariant; indeed if $\alpha\in R$, then $x\alpha=\alpha$, and hence for every $y\in C_{\Gamma}(x)$ we have $xy\alpha=yx\alpha=y\alpha$, so $y\alpha\in R$. The extension $R/F$ is algebraic, so $R$ is the real closure of $F$. Therefore, the Galois group $\mathrm{Gal}(R/F)$ is trivial by [18, Theorem 2.9, page 455]. We get that $C_{\Gamma}(x)$ lies in the subgroup of $\Gamma$ fixing $R$, so by the profinite version of the Galois correspondence $C_{\Gamma}(x)\subseteq C$. ∎
This shows that Theorem 2.1 has the following consequence:
Corollary 2.4.
The absolute Galois group of a field $F$ is real projective if and only if $F$ satisfies $\mathrm{cd}(F(\mathbf{i}))\leq 1$. ∎
Definition 2.5.
Let $F$ be a field. The real spectrum $\mathrm{Spr}(F)$ of $F$ is the set of all orderings of $F$. We consider $\mathrm{Spr}(F)$ as a topological space by equipping it with the Harrison topology. The latter has the following explicit description: let $S\subset F$ be a finite subset and let $\prec$ be a complete ordering on $S$. Let $U(S,\prec)\subset\mathrm{Spr}(F)$ be the set consisting of all orderings $<$ on $F$ such that $<$ restricted to $S$ is $\prec$. The sets $U(S,\prec)$ form a sub-basis of the Harrison topology, which makes $\mathrm{Spr}(F)$ a compact and totally disconnected topological space.
Definition 2.6.
For every $<\in\mathrm{Spr}(F)$, let $F_{<}$ denote the real closure of the ordered field $(F,<)$. We say that the field $F$ is pseudo real closed if every absolutely irreducible variety defined over $F$ which has a simple $F_{<}$-rational point for every $<\in\mathrm{Spr}(F)$ has an $F$-rational point.
Remark 2.7.
Recall that a field $F$ is pseudo algebraically closed if every geometrically irreducible variety defined over $F$ has a $F$-rational point. By a theorem of Ax [2], the absolute Galois group $\Gamma$ of such a field is projective, and in particular it has cohomological dimension at most one. By Prestel’s Extension Theorem [31, Theorem 3.1, page 148], every algebraic extension of a pseudo real closed field is pseudo real closed. This implies that if $F$ is pseudo real closed, then $F(\mathbf{i})$ is pseudo algebraically closed since $F(\mathbf{i})$ has no orderings. Hence $F$ has virtual cohomological dimension $\leq 1$, and in particular $\Gamma(F)$ is real projective by Haran’s theorem.
The following result of Haran–Jarden is a strong converse of the remark above (see [11, Theorem 10.4 on page 487]).
Theorem 2.8 (Haran–Jarden).
For every real projective profinite group $G$ there is a pseudo real closed field $F$ whose absolute Galois group is isomorphic to $G$. ∎
Remark 2.9.
Therefore the Hopkins–Wickelgren formality conjecture for the class of fields of virtual cohomological dimension $\leq 1$ is the same as for the class of pseudo real closed fields, and it is a purely group-theoretical problem for the class of real projective profinite groups.
Definition 2.10.
Let $\Gamma$ be a profinite group. Let $\mathcal{Y}(\Gamma)$ denote the subset of elements of order $2$ in $\Gamma$. We equip $\mathcal{Y}(\Gamma)$ with the subset topology. Let $\mathcal{X}(\Gamma)$ denote the quotient of $\mathcal{Y}(\Gamma)$ by the conjugation action of $\Gamma$. We equip $\mathcal{X}(\Gamma)$ with the quotient topology. Note that when $\Gamma$ is the absolute Galois group of a field $F$ then $\mathcal{X}(\Gamma)$ is just the real spectrum of $F$ equipped with its Harrison topology by classical Artin–Schreier theory.
Let $C(\mathcal{X}(\Gamma),\mathbb{F}_{2})$ be the ring of continuous functions from $\mathcal{X}(\Gamma)$ to $\mathbb{\mathbb{F}}_{2}$ where we equip the latter with the discrete topology.
Theorem 2.11 (Scheiderer).
Let $\Gamma$ be a real projective group and $\mathcal{X}(\Gamma)$ be as above. Then, for each $n\geq 0$, there is a natural homomorphism
(2) $\pi_{n}\colon H^{n}(\Gamma,\mathbb{F}_{2})\to C(\mathcal{X}(\Gamma),\mathbb{F}_{2})$
which is an isomorphism for $n>1$ and surjective for $n=1$.
Proof.
In [33] Scheiderer constructs his theory of real étale cohomology in two cases: for schemes over which $2$ is invertible, and for profinite groups. In the case we are interested in, real projective groups, the two approaches are actually equivalent by Theorem 2.8. The assertion then follows from [34, Proposition 2.15].
We will briefly indicate how to use Theorem 2.8 and the results of [33] to deduce the result. By Theorem 2.8 we can find a field $F$ such that $\Gamma=\Gamma(F)$. By [33, Theorem 6.6, page 61] applied to the spectrum of $F$ and the constant discrete $\Gamma$-module $A=\mathbb{F}_{2}$, there is a long exact sequence of the form
(3) $\cdots\to H^{n}_{b}(F,\mathbb{F}_{2})\to H^{n}(\Gamma,\mathbb{F}_{2})\xrightarrow{p^{n}_{F}}H^{0}(\mathcal{X}(\Gamma),\mathcal{H}^{n}(\mathbb{F}_{2}))\to H^{n+1}_{b}(F,\mathbb{F}_{2})\to\cdots$
where $\mathcal{H}^{n}(\mathbb{F}_{2})$ denotes the $\Gamma$-equivariant higher direct image sheaf constructed in [33, Section 8] (see especially [33, Remark 8.9, page 95]), and $H_{b}$ denotes cohomology group with respect to the $b$-topology on $\mathrm{Spec}\,F$ as defined in [33, Definition 2.3, page 11]. For the identification of sequence (3) with the long exact sequence of [33, Theorem 6.6, page 61], we recall that $\mathcal{X}(\Gamma)$ is the real spectrum of the field $F$ (see [33, 9.2, page 97]). We can then use [33, Remarks 9.5 and 9.6, page 99] to identify the remaining terms. In particular, we can identify the map $p^{n}_{F}$ in (3) with the map $\pi_{n}$ in (2). By [33, Theorem 7.3, page 70] (see also [33, Corollary 9.8]), the cohomological $2$-dimension of $F$ with respect to the $b$-topology is equal to the virtual cohomological $2$-dimension of $\Gamma$. Since the latter is $1$ by Theorem 2.1 and Corollary 2.4, the group $H^{n}_{b}(F,\mathbb{F}_{2})$ in sequence (3) vanishes for $n>1$. Hence, since sequence (3) is exact, the map $\pi_{n}$ in (2) is an isomorphism for $n>1$ and is still surjective for $n=1$. ∎
Remark 2.12.
We note that there is also a group-theoretic proof of Theorem 2.11 in [33, Section 12]. Since the virtual cohomological $2$-dimension of a real projective group is $1$, and $\Gamma$ contains no subgroup isomorphic to $\mathbb{Z}/2\times\mathbb{Z}/2$ by Haran’s Theorem 2.1, [33, Theorem 12.13, page 151] applies with $A=\mathbb{F}_{2}$, and the map $\pi_{n}$ in (2) is an isomorphism for every $n>1$. The case $n=1$ is proven in [33, Remark 12.20.2, page 158] where the argument of the second paragraph applies since every involution is self-centralising. A proof of Theorem 2.11 using Artin–Schreier structures which are discussed in [33, Appendix B] can also be deduced from [33, Proposition B.8.1, page 257].
3. Massey products and formality
In this section we show that formality implies strong Massey vanishing. This is a well known fact which can for example be deduced from the theory of $A_{\infty}$-algebras. However, for a lack of reference known to the authors to a proof which avoids the $A_{\infty}$-structures, especially in positive characteristic, and for the convenience of the reader we provide direct and self-contained proofs for the assertions needed. In this section we only assume familiarity with some basic terminology for model categories, and we recall and use a basic fact about weak equivalences in Lemma 3.4.
Notation 3.1.
Let $\mathbb{F}$ be a field, and let $\mathbf{DGA}$ denote the category of associative differential graded algebras over $\mathbb{F}$. We will often refer to the objects in $\mathbf{DGA}$ just as dg-algebras. We will often denote the multiplication on objects in $\mathbf{DGA}$ as a cup product $\cup$. Let $C^{*}$ be a dg-algebra with differential $\delta\colon C^{*}\to C^{*+1}$, and cohomology $H^{*}=\textrm{Ker}(\delta)/\textrm{Im}(\delta)$. We will often write $C$ instead of $C^{*}$.
We consider $\mathbf{DGA}$ as a model category with the standard model structure, as constructed in [14, Theorem 5, page 58]. It is determined by the following classes of weak equivalences and fibrations: a map is a weak equivalence if its underlying map of chain complexes is a quasi-isomorphism; and it is a fibration if it is degree-wise surjective. Recall that maps which are both weak equivalences and fibrations are also called acyclic fibrations. The cofibrations are the maps which have the left lifting property with respect to all acyclic fibrations.
Definition 3.2.
Let $C$ be a dg-algebra with cohomology $H$. We consider $H$ as an object in $\mathbf{DGA}$ with trivial differential. Then $C$ is called formal if there is a diagram
(4) $C\leftarrow T\to H$
of weak equivalences of dg-algebras.
Remark 3.3.
We recall that, since all objects are fibrant in the standard model structure on $\mathbf{DGA}$, a sequence of quasi-isomorphisms
$C\leftarrow T_{1}\to T_{2}\leftarrow T_{3}\to\ldots\leftarrow T_{n}\to B$
in $\mathbf{DGA}$ can be reduced to a sequence $C\leftarrow T\to B$ of just two quasi-isomorphism.
Lemma 3.4.
Let $f\colon A\to B$ be a weak equivalence of dg-algebras. Then there is a commutative diagram of morphisms of dg-algebras
such that all maps are weak equivalences, the map $j$ is a homotopy equivalence, and $p$ and $q$ are fibrations.
Proof.
This is a formal consequence of the model structure on $\mathbf{DGA}_{\mathbb{F}}$. Let
[…]
Since $\widetilde{A}$ is cofibrant and $q$ is an acyclic fibration, there is a dotted lift $j$ which makes the diagram commute. As $f\circ p$ is also a weak equivalence, $j$ must be a weak equivalence. Since $\widetilde{A}$ and $\widetilde{B}$ are cofibrant and fibrant objects (every object in $\mathbf{DGA}$ is fibrant), $j$ is a homotopy equivalence, in the sense that there is a map $\widetilde{B}\xrightarrow{\widetilde{h}}\widetilde{A}$ which is an inverse to $j$ up to homotopy. This proves the claim. ∎
The existence of diagram (9) will allow us to lift Massey products also along quasi-isomorphisms of dg-algebras.
Proposition 3.5.
Let $f\colon A\to B$ be a weak equivalence of dg-algebras, and let $a_{1},a_{2},\ldots,a_{n}$ be classes in $H^{*}(A)$. Let $b_{1},b_{2},\ldots,b_{n}$ be their images in $H^{*}(B)$. Then $f_{*}\colon H^{*}(A)\to H^{*}(B)$ maps the Massey product set $\langle a_{1},a_{2},\ldots,a_{n}\rangle$ bijectively onto the Massey product set $\langle b_{1},b_{2},\ldots,b_{n}\rangle$.
Proof.
Let $\{a_{i,i+j}\}\subset A$ for $11\leq i<i+j\leq n+$ 1 be a defining system for the $n$-fold Massey product of $a_{1},\ldots,a_{n}$. Since $f$ commutes with the differentials and cup-products, it sends the defining system $\{a_{i,i+j}\}$ to a defining system $\{b_{i,i+j}:=f(a_{i,i+j})\}\subset B$ for the $n$-fold Massey product of $b_{1},\ldots,b_{n}$. Hence $f_{*}$ sends elements in the Massey product set $\langle a_{1},a_{2},\ldots,a_{n}\rangle$ to elements in the Massey product set $\langle b_{1},b_{2},\ldots,b_{n}\rangle$. More concretely, $f_{*}$ sends the class of the cocycle $\sum_{k=2}^{n}\bar{a}_{1,k}\cup a_{k,n+1}$ in $\langle a_{1},\ldots,a_{n}\rangle$ to the class of the cocycle $\sum_{k=2}^{n}\bar{b}_{1,k}\cup b_{k,n+1}$ in $\langle b_{1},\ldots,b_{n}\rangle$. Since $f_{*}$ is an isomorphism and therefore injective, it remains to show that every defining system for the $n$-fold Massey product of $b_{1},\ldots,b_{n}$ can be lifted to a defining system for the $n$-fold Massey product of $a_{1},\ldots,a_{n}$.
By Lemma 3.4, in order to prove the claim for $f$, it suffices to prove the corresponding claims for the maps $p$, $q$ and $j$ as in diagram (9). The claim for $j$ follows from the following lemma.
Lemma 3.6.
If $f$ is a homotopy equivalence of dg-algebras, then $f_{*}$ restricts to a bijection between $\langle a_{1},a_{2},\ldots,a_{n}\rangle$ and $\langle b_{1},b_{2},\ldots,b_{n}\rangle$.
Proof.
Since $f$ is a homotopy equivalence, we can choose a morphism of dg-algebras $g\colon B\to A$ of $f$ such that there are quasi-isomorphisms $g\circ f\simeq\mathrm{id}_{A}$ and $f\circ g\simeq\mathrm{id}_{B}$. In particular, $g$ is itself a quasi-isomorphism of dg-algebras, and $g$ sends the defining system $\{b_{i,i+j}\}$ for the $n$-fold Massey product of $b_{1},\ldots,b_{n}$ to a defining system $\{a^{\prime}_{i,i+j}:=g(b_{i,i+j})\}\subset A$ for the $n$-fold Massey product of $a_{1},\ldots,a_{n}$. The defining system $\{a^{\prime}_{i,i+j}:=g(b_{i,i+j})\}$ is a lift of the defining system $\{b_{i,i+j}\}$. In fact, since $\sum_{k=2}^{n}\bar{b}_{1,k}\cup b_{k,n+1}$ is a cocycle in $A$, the quasi-isomorphism $g\circ f\simeq\mathrm{id}_{A}$ implies that
$\sum_{k=2}^{n}\bar{a}_{1,k}\cup a_{k,n+1}-\sum_{k=2}^{n}\bar{a}^{\prime}_{1,k}\cup a^{\prime}_{k,n+1}$
is a coboundary in $A$. Thus, the cohomology classes of the cocycles $\sum_{k=2}^{n}\bar{a}_{1,k}\cup a_{k,n+1}$ and $\sum_{k=2}^{n}\bar{a}^{\prime}_{1,k}\cup a^{\prime}_{k,n+1}$, respectively, are the same element in the Massey product set $\langle a_{1},a_{2},\ldots,a_{n}\rangle\subset H^{*}(A)$. ∎
To prove the claim of Proposition 3.5 for the acyclic fibrations $p$ and $q$, we are going to use the following basic statements about quasi-isomorphisms.
Lemma 3.7.
Let $f\colon A\to B$ be an acyclic fibration of dg-algebras. Let $b\in B^{n}$ be such that $\delta b=0$. Then there is an $a\in A^{n}$ such that $f(a)=b$ and $\delta a=0$.
Proof.
Since $f$ is a quasi-isomorphism, there is an $a^{\prime}\in A^{n}$ such that $\delta a^{\prime}=0$ and $f(a^{\prime})$ and $b$ have the same cohomology class. In other words, there is a $c\in B^{n-1}$ such that $f(a^{\prime})=b-\delta c$. Since $f$ is surjective, there is a $d\in A^{n-1}$ such that $f(d)=c$. Set $a=a^{\prime}+\delta d$. Then $\delta a=\delta a^{\prime}+\delta^{2}d=0$, and
$f(a)=f(a^{\prime})+f(\delta d)=f(a^{\prime})+\delta f(d)=b-\delta c+\delta c=b.\qed$
Lemma 3.8.
Let $f\colon A\to B$ be an acyclic fibration of dg-algebras. Let $d\in A^{n}$ and $c\in B^{n-1}$ be such that $\delta d=0$ and $\delta c=f(d)$. Then there is a $\bar{d}\in A^{n-1}$ such that $f(\bar{d})=c$ and $\delta\bar{d}=d$.
Proof.
Since $f$ is a quasi-isomorphism and the cohomology class of $f(d)$ is trivial, there is a $d^{\prime}\in A^{n-1}$ such that $\delta d^{\prime}=d$. Then $\delta c=f(\delta d^{\prime})=\delta f(d^{\prime})$, so $\delta(c-f(d^{\prime}))=0$. By Lemma 3.7 above there is an $e\in A^{n-1}$ such that $\delta e=0$ and $f(e)=c-f(d^{\prime})$. Set $\bar{d}=d^{\prime}+e$. Then $\delta(\bar{d})=\delta(d^{\prime})+\delta(e)=d+0=d$, and $f(\bar{d})=f(d^{\prime})+f(e)=f(d^{\prime})+c-f(d^{\prime})=c$. ∎
We are now ready to conclude the proof of Proposition 3.5. Since $p$ and $q$ are acyclic fibrations, the proof is reduced to the case when $f$ is an acyclic fibration. Therefore we assume from now on that $f$ is an acyclic fibration. We need to show that any defining system for the $n$-fold Massey product of $b_{1},\ldots,b_{n}$ can be lifted to a defining system of $a_{1},\ldots,a_{n}$. Let $\{b_{i,i+j}\}\subset B$ for $11\leq i<i+j\leq n+$ 1 be a defining system for the $n$-fold Massey product of $b_{1},\ldots,b_{n}$. By Lemma 3.7, we can choose cocycles $a_{1,2},a_{2,3},\ldots,a_{n,n+1}\in A$ with $f(a_{i,i+1})=b_{i,i+1}$ and $a_{i,i+1}$ represents the cohomology class $a_{i}$ for each $i$. Next, for a given $j\geq 2$, assume that we have constructed elements $a_{i,i+l}$ for all $i$ and all $l<j$ satisfying $f(a_{i,i+l})=b_{i,i+l}$ and
(10) $\delta a_{i,i+l}$ $=\sum_{k=i+1}^{i+l-1}\bar{a}_{i,k}\cup a_{k,i+l}$
where we recall that $d_{i,k}$ denotes the degree of $a_{i,k}$ and $\bar{a}_{i,k}=(-1)^{1+d_{i,k}}a_{i,k}$. We claim that $d:=\sum_{k=i+1}^{i+j-1}\bar{a}_{i,k}\cup a_{k,i+j}$ is a cocycle. Applying $\delta$ and using $(-1)^{1+d_{i,k}}\cdot(-1)^{d_{i,k}}=-1$ yields
$\delta(d)=\sum_{k=i+1}^{i+j-1}(-1)^{1+d_{i,k}}\delta(a_{i,k})\cup a_{k,i+j}-a_{i,k}\cup\delta(a_{k,i+j}).$
Using $\delta(a_{i,i+1})=0=\delta(a_{i+j-1,i+j})$ and the relations given by (10) we get
$\delta(d)=$ $\sum_{k=i+2}^{i+j-1}(-1)^{1+d_{i,k}}\left(\sum_{s=i+1}^{k-1}(-1)^{1+d_{i,s}}a_{i,s}\cup a_{s,k}\right)\cup a_{k,i+j}$
$-\sum_{k=i+1}^{i+j-2}a_{i,k}\cup\left(\sum_{r=k+1}^{i+j-1}(-1)^{1+d_{kr}}a_{k,r}\cup a_{r,i+j}\right).$
Now we use that, for each triple $i<s<k$, we have $1+d_{ik}=d_{is}+d_{sk}$ and get
$\delta(d)=\sum_{k=i+2}^{i+j-1}\left(\sum_{s=i+1}^{k-1}a_{i,s}\cup\bar{a}_{s,k}\right)\cup a_{k,i+j}-\sum_{k=i+1}^{i+j-2}a_{i,k}\cup\left(\sum_{r=k+1}^{i+j-1}\bar{a}_{k,r}\cup a_{r,i+j}\right)$
$=$ $\left(\sum_{i+1\leq s<k\leq i+j-1}a_{i,s}\cup\bar{a}_{s,k}\cup a_{k,i+j}\right)-\left(\sum_{i+1\leq k<r\leq i+j-1}a_{i,k}\cup\bar{a}_{k,r}\cup a_{r,i+j}\right)=0.$
Hence we can apply Lemma 3.8 with $d=\sum_{k=i+1}^{i+j-1}(-1)^{1+d_{i,k}}a_{i,k}\cup a_{k,i+j}$ and $c=b_{i,i+j}$ to get an element $a_{i,i+j}\in A$ which satisfies
$f(a_{i,i+j})=b_{i,i+j}~{}\text{and}~{}\delta a_{i,i+j}=\sum_{k=i+1}^{i+j-1}\bar{a}_{i,k}\cup a_{k,i+j}.$
Continuing this process we get a defining system for the classes $a_{1},a_{2},\ldots,a_{n}$. ∎
As a consequence we get that formality implies strong Massey vanishing:
Theorem 3.9.
Every formal differential graded algebra satisfies strong Massey vanishing.
Proof.
Let $C$ be a formal dg-algebra and let $H$ denote the cohomology of $C$. By Proposition 3.5, it suffices to show that if we consider $H$ as a dg-algebra with trivial differential, then for every $n$-tuple $a_{1},\ldots,a_{n}$ in $H$ of elements whose neighbouring cup products vanish the corresponding Massey product is defined and vanishes in $H$. Indeed, let $\{a_{i,i+j}\}\in H$ for $11\leq i<j\leq n+$ 1 and $(i,j)\neq(1,n+1)$ be such that $a_{i,i+1}=a_{i}$ for every index $i$ and $a_{i,j}=0$, if $|i-j|>1$. Then $\{a_{i,i+j}\}$ is a defining system for the Massey product of $a_{1},\ldots,a_{n}$ whose corresponding Massey product vanishes. ∎
4. Graded Hochschild cohomology and formality
One way to prove the formality of a dg-algebra is to show that certain Hochschild cohomology groups vanish. We are first going to recall the definition of graded Hochschild cohomology and then discuss the relation to formality. Since we only consider algebras and dg-algebras over the field $\mathbb{F}_{2}$, we will from now on work over $\mathbb{F}_{2}$. Unless stated otherwise $\otimes$ will denote $\otimes_{\mathbb{F}_{2}}$.
Definition 4.1.
Let $A=\bigoplus_{i\geq 0}A_{i}$ be a graded-commutative $\mathbb{F}_{2}$-algebra. A graded $A$-module is an $A$-module $M$ with a decomposition $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ into subgroups such that $A_{i}\cdot M_{j}\subseteq M_{i+j}$ for every $i\geq 0,j\in\mathbb{Z}$. A homomorphism of graded $A$-modules from $M=\bigoplus_{i\in\mathbb{Z}}M_{i}$ to $N=\bigoplus_{i\in\mathbb{Z}}N_{i}$ is an $A$-module homomorphism $\phi\colon M\to N$ which respects the gradings, i.e., we have $\phi(M_{i})\subseteq N_{i}$ for every $i$. Let $\underline{\mathrm{Hom}}_{A}(M,N)$ denote the set of graded $A$-modules from $M$ to $N$; it has the structure of an $A$-module. Graded $A$-modules with graded $A$-module homomorphisms form an abelian category.
Let $A^{e}$ denote the tensor product $A\otimes A$. It has a structure of a commutative graded $\mathbb{F}_{2}$-algebra with respect to the decomposition:
$A^{e}=\bigoplus_{n\geq 0}\bigoplus_{i+j=n}A_{i}\otimes A_{j}.$
Note that every graded $A$-module $M$ has the structure of a graded $A^{e}$-module: we equip $M$ with the unique $A^{e}$-module structure such that for every $a,a^{\prime}\in A$ and $m\in M$ we have
$(a\otimes a^{\prime})m=aa^{\prime}m,$
and we keep its grading.
For $s\in\mathbb{Z}$ and a graded $A$-module $M$, we write $M[s]$ for the graded $A$-module given in degree $n$ by $M[s]_{n}=M_{n+s}$. It is straight-forward to show the following:
Lemma 4.2.
We have
$\mathrm{Hom}_{A^{e}}(A,M)=\bigoplus_{s\in\mathbb{Z}}\underline{\mathrm{Hom}}_{A^{e}}(A,M[s])$
for every graded $A^{e}$-module $M$, where $\mathrm{Hom}_{A^{e}}(\cdot,\cdot)$ denotes the module of all $A^{e}$-module homomorphisms, ignoring the gradings. ∎
We now define the graded Hochschild cohomology groups of an $\mathbb{F}_{2}$-algebra $A$ with coefficients in a graded $A^{e}$-module $M$. First, we recall the bar complex (see e.g., [40, Section 1.1]). For $n\geq 0$, we write $B_{n}(A)=A^{\otimes(n+2)}$ and define the map $(d_{B})_{n}=d_{n}\colon B_{n}(A)\to B_{n-1}(A)$ by
$d_{n}(a_{0}\otimes a_{1}\otimes\cdots\otimes a_{n+1})=\sum_{i\geq 0}^{n}a_{0}\otimes\cdots\otimes a_{i}a_{i+1}\otimes\cdots\otimes a_{n+1}$
for $a_{0},\ldots,a_{n+1}\in A$. Then the bar complex of the $A^{e}$-module $A$ is the complex
[…]
where $\mu$ denotes the multiplication on $A$. For every $n\geq 0$, $B_{n}(A)$ may be considered as an $A^{e}$-module via
$(a\otimes a^{\prime})\cdot(a_{0}\otimes\cdots\otimes a_{n+1})=(aa_{0})\otimes a_{1}\otimes\cdots\otimes a_{n}\otimes(a_{n+1}a^{\prime}).$
We equip $B_{n}(A)$ with the grading induced by the grading of $A$, and we consider $d_{n}$ as a graded homomorphism.
Definition 4.3.
Let $M$ be a graded $A^{e}$-module. The graded Hochschild cohomology group $\mathrm{HH}^{n,s}(A,M)$ is defined as the $n$th cohomology of the cochain complex $(\underline{\mathrm{Hom}}_{A^{e}}(B_{*}(A),M[s]),d_{B}^{*})$ with differential $d_{B}^{*}f=f\circ d_{B}$.
Remark 4.4.
We now recall that graded Hochschild cohomology can be computed using a slightly simpler complex. For every $n$, there is a natural $\mathbb{F}_{2}$-linear isomorphism
(11) $\underline{\mathrm{Hom}}_{A^{e}}(B_{n}(A),M)\xrightarrow{\cong}\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes n},M)$
defined by sending a map $f$ to the $\mathbb{F}_{2}$-linear map which sends $a_{1}\otimes\cdots\otimes a_{n}$ to $f(1\otimes a_{1}\otimes\cdots\otimes a_{n}\otimes 1)$. The maps $d^{*}_{n}\colon\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes n},M)\to\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes(n+1)},M)$, defined for every $n$ by
$d^{*}_{n}(f)(a_{1}\otimes\cdots\otimes a_{n+1})$ $=a_{1}f(a_{2}\otimes\cdots\otimes a_{n+1})$
$+\sum_{i=1}^{n}f(a_{1}\otimes\cdots\otimes a_{i}a_{i+1}\otimes\cdots\otimes a_{n+1})$
$+f(a_{1}\otimes\cdots\otimes a_{n})a_{n+1},$
turn $(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes*},M),d^{*})$ into a cochain complex. We can then check that isomorphism (11) induces an isomorphism of cochain complexes
$(\underline{\mathrm{Hom}}_{A^{e}}(B_{*}(A),M),d_{B}^{*})\xrightarrow{\cong}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes*},M),d^{*}).$
Hence, for every $n$ and $s$, we have an isomorphism
(12) $\mathrm{HH}^{n,s}(A,M)\cong H^{n}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes*},M[s]),d^{*}).$
In fact, isomorphism (12) is used to define Hochschild cohomology in [35, Section 4, page 85] (see also [3, Section 4]).
We will now recall a criterion on the Hochschild cohomology of a graded algebra which implies formality. In fact, it implies the following stronger property.
Definition 4.5.
A graded $\mathbb{F}_{2}$-algebra $A$ is called intrinsically formal if every differential graded algebra $C$ with $H^{*}(C)\cong A$ is formal.
Remark 4.6.
We note that intrinsic formality is a property formulated for graded algebras while formality is a property formulated for dg-algebras. However, one may call a dg-algebra $C$ intrinsically formal if its cohomology algebra is. Intrinsic formality is a much stronger property than formality.
An equivalent characterisation of intrinsic formality is given as follows:
Lemma 4.7.
Let $A$ be a graded $\mathbb{F}_{2}$-algebra. Then $A$ is intrinsically formal if and only if any two differential graded algebras with cohomology isomorphic to $A$ are quasi-isomorphic.
Proof.
Let $C$ be a dg-algebra with $H^{*}(C)\cong A$. If any two dg-algebras with cohomology isomorphic to $A$ are quasi-isomorphic, then $C$ and $H^{*}(C)$ are quasi-isomorphic. Hence $C$ is formal. This shows that $A$ is intrinsically formal.
To prove the other implication, let $C_{1}$ and $C_{2}$ be dg-algebras with cohomology algebras $H^{*}(C_{1})\cong A\cong H^{*}(C_{2})$. If $A$ is intrinsically formal, then $C_{1}$ and $C_{2}$ are formal. This means that there are quasi-isomorphisms of dg-algebras $C_{1}\leftarrow T_{1}\to A$ and $A\leftarrow T_{2}\to C_{2}$ where we consider $A$ as a dg-algebra with trivial differential. By Remark 3.3, we can then find a dg-algebra $T$ and quasi-isomorphisms of dg-algebras $C_{1}\leftarrow T\to C_{2}$. Thus, $C_{1}$ and $C_{2}$ are quasi-isomorphic. ∎
The following theorem is due to Kadeishvili (see [16] and [17]). For a proof we refer to the work of Seidel and Thomas [35, Theorem 4.7, page 85].
Theorem 4.8 (Kadeishvili).
Let $A=\bigoplus_{i\geq 0}A_{i}$ be a non-negatively graded $\mathbb{F}_{2}$-algebra with $A_{0}=\mathbb{F}_{2}$. Assume that
$\mathrm{HH}^{n,2-n}(A)=0~{}\text{for all}~{}n\geq 3.$
Then $A$ is intrinsically formal. ∎
Theorem 4.8 and Lemma 4.7 then imply:
Corollary 4.9.
Let $C$ be a dg-algebra over $\mathbb{F}_{2}$ with $C^{0}=\mathbb{F}_{2}$. Assume that
$\mathrm{HH}^{n,2-n}(H^{*}(C))=0~{}\text{for all}~{}n\geq 3.$
Then $C$ is formal. ∎
We will refer to the above criteria using the following terminology:
Definition 4.10.
For a non-negatively graded $\mathbb{F}_{2}$-algebra $A=\bigoplus_{i\geq 0}A_{i}$ with $A_{0}=\mathbb{F}_{2}$, we say that $A$ satisfies Kadeishvili’s vanishing if
$\mathrm{HH}^{n,2-n}(A)=0\textrm{ for every $n\geq 3$.}$
For a dg-algebra $C$ over $\mathbb{F}_{2}$ with $C^{0}=\mathbb{F}_{2}$, we say that $C$ satisfies Kadeishvili’s vanishing if the graded $\mathbb{F}_{2}$-algebra $H^{*}(C)$ satisfies Kadeishvili’s vanishing.
Remark 4.11.
For the reader familiar with $A_{\infty}$-algebras we note that the condition of Theorem 4.8 can be interpreted and the proof can be summarised as follows: Let $\mathcal{A}$ be an $A_{\infty}$-algebra and, for $n\geq 1$, let $m_{n}\colon\mathcal{A}^{\otimes n}\to\mathcal{A}$ denote the maps of degree $2-n$ which define the $A_{\infty}$-structure. Then each $m_{n}$ is a cocycle in Hochschild cohomology of $H^{*}(\mathcal{A})$ which represents a class in $\mathrm{HH}^{n,2-n}(H^{*}(\mathcal{A}))=0$. Now let $A$ be a graded algebra and let $C$ be a dg-algebra with an isomorphism of graded algebras $A\xrightarrow{\cong}H^{*}(C)$. Recall that we can view $A$ and $C$ as $A_{\infty}$-algebras with trivial higher multiplication maps $m_{n}$ for $n\geq 3$. Kadeishvili showed that it is always possible to equip $A$ with the structure of an $A_{\infty}$-algebra such that there is a quasi-isomorphism of $A_{\infty}$-algebras $q\colon A\to C$ which lifts the isomorphism $A\xrightarrow{\cong}H^{*}(C)$ of graded algebras (see [16] and [17]). Since the higher level $A_{\infty}$-maps $m_{n}$ on $A$ may be nontrivial for $n\geq 3$, the $A_{\infty}$-algebra structure on $A$ is different from the one it inherits as a graded algebra. Moreover, $q$ is not morphism of dg-algebras in general. However, if $\mathrm{HH}^{n,2-n}(A)=0$ for all $n\geq 3$, then all higher multiplication maps on $A$ must be trivial. One can then lift the $A_{\infty}$-morphism $q\colon A\to C$ via a certain tensor algebra $T$ to a zigzag of quasi-isomorphisms $A\leftarrow T\to C$ of dg-algebras. Then it suffices to compose with the algebra isomorphism $H^{*}(C)\cong A$ to conclude that $C$ is formal. We refer to [35, Proof of Theorem 4.7] for the details.
5. Graded Hochschild cohomology of Koszul algebras
We continue to work over the field $\mathbb{F}_{2}$. Unless stated otherwise $\otimes$ will denote $\otimes_{\mathbb{F}_{2}}$.
Definition 5.1.
A quadratic algebra is a non-negatively graded $\mathbb{F}$-algebra $A=\bigoplus_{i\geq 0}A_{i}$ such that $A_{0}=\mathbb{F}$ and $A$ is generated over $\mathbb{F}_{2}$ by $A_{1}$ with relations of degree two. Explicitly, let
$T(A_{1})=\mathbb{F}_{2}\oplus A_{1}\oplus(A_{1}\otimes A_{1})\oplus\cdots=\bigoplus_{i\geq 0}A_{1}^{\otimes i}$
be the free tensor algebra of the $\mathbb{F}_{2}$-vector space $A_{1}$. Then $A$ is quadratic if the canonical map $\tau\colon T(A_{1})\to A$ is surjective and $\ker(\tau)$ is generated by its component $R=\ker(\tau)\cap(A_{1}\otimes A_{1})$ of degree two, so that $A=T(A_{1})/(R)$, where $(R)$ denotes the ideal generated by $R$. A quadratic algebra $A$ is called locally finite if each $A_{i}$ is a finite-dimensional vector space over $\mathbb{F}$.
For a quadratic algebra $A=\bigoplus_{i\geq 0}A_{i}$ we denote by $A_{+}$ the ideal $A_{+}=\bigoplus_{i>0}A_{i}$. We will later use the following natural way to create new quadratic algebras out of old (see for e.g. [27, Chapter 3.1]):
Definition 5.2.
Let $A=\bigoplus_{i\geq 0}A_{i}$ and $B=\bigoplus_{i\geq 0}B_{i}$ be quadratic algebras. The connected sum of $A$ and $B$, denoted by $A\sqcap B$, is the graded ring with $(A\sqcap B)_{0}=\mathbb{F}_{2}$, $(A\sqcap B)_{i}=A_{i}\oplus B_{i}$ for $i>0$ and multiplication $A_{+}B_{+}$ and $B_{+}A_{+}$ is set to be zero.
The connected sum $C:=A\sqcap B$ is a quadratic algebra. Let $R_{A}\subseteq A_{1}\otimes A_{1}$ and $R_{B}\subseteq B_{1}\otimes B_{1}$ denote the relations in the tensor algebras defining $A$ and $B$, respectively. Then we can write $C$ as a quotient of a tensor algebra
$C=T(A_{1}\oplus B_{1})/(R_{C})$
where $R_{C}\subseteq C_{1}^{\otimes 2}$ is given by $R_{C}=R_{A}\oplus R_{B}\oplus(A_{1}\otimes B_{1})\oplus(B_{1}\otimes A_{1})$.
To every quadratic algebra we can associate the following chain complex:
Definition 5.3.
Let $V$ be an $\mathbb{F}_{2}$-vector space and let $T(V)=\bigoplus_{i\geq 0}V^{\otimes i}$ be its tensor algebra. We consider $T(V)$ as a graded $\mathbb{F}_{2}$-algebra with $|v|=1$ for all $v\in V$. Let $R\subset V\otimes V$ and let $A=T(V)/(R)$ be the associated quadratic algebra where $(R)$ denotes the ideal generated by $R$ in $T(V)$. We denote by $K_{i}^{i}(A)$ the $\mathbb{F}_{2}$-linear subspace defined by $K_{0}^{0}(A)=\mathbb{F}_{2}$, $K_{1}^{1}(A)=V$, $K_{2}^{2}=R$ and
$K_{i}^{i}(A)=\bigcap_{0\leq j\leq i-2}V^{\otimes j}\otimes R\otimes V^{\otimes i-j-2}\subset V^{\otimes i}~{}\text{for}~{}i\geq 3.$
We set $K_{i}(A)=A\otimes K_{i}^{i}(A)\otimes A$. We consider $K_{i}(A)$ as an $A^{e}$-module with the $A$-bimodule structure induced by the left- and right action of $A$ on the leftmost and rightmost factor $A$, respectively. For each $i\geq 0$, we define a homomorphism $d=d_{i}\colon K_{i}(A)\to K_{i-1}(A)$ by
$d_{i}(a\otimes x_{1}\otimes\cdots\otimes x_{i}\otimes a^{\prime})=(ax_{1})\otimes x_{2}\otimes\cdots\otimes x_{i}\otimes a^{\prime}+a\otimes x_{1}\otimes x_{2}\otimes\cdots\otimes(x_{i}a^{\prime}).$
Since $R\subset V^{\otimes 2}$ generates the relations in $A$, it is clear that $d^{2}=0$. We refer to the chain complex $(K(A),d)$ as the Koszul complex of $A$.
Let $(B(A),d_{B})$ denote the bar complex of $A$ (see e.g. in [40, (1.1.4)]). Since $R$ describes the relations in $A$, the natural inclusion $K_{i}(A)\hookrightarrow B_{i}(A)$ defines a morphism of complexes $(K(A),d)\to(B(A),d_{B})$. The multiplication map $\mu\colon A\otimes A\to A$ defines morphisms of complexes from the complexes $(K(A),d)$ and $(B(A),d_{B})$ to the complex $A$ concentrated in degree 00 with trivial differential. However, while $\mu\colon B(A)\to A$ is always a quasi-isomorphism, $\mu\colon K(A)\to A$ may not be a quasi-isomorphism for an arbitrary quadratic algebra $A$. This leads to the following definition.
Definition 5.4.
A quadratic algebra $A$ is called Koszul if the morphism of complexes induced by multiplication $\mu\colon K(A)\to A$ is a quasi-isomorphism.
We note that there are many equivalent definitions of the notion of a Koszul algebra. By [38, Proposition 3.1](where the notation $K^{\prime}(A)$ is used for $K(A)$), Definition 5.4 is equivalent to the ones in [27], [29] and [30] (see also [4, Section 2.1]). Again, by [38, Proposition 3.1], the Koszul complex $(K(A),d)$ is a complex of free $A^{e}$-modules. Hence, if $A$ is a Koszul algebra, we can use $K(A)$ to compute Hochschild cohomology. In fact, since $K(A)\hookrightarrow B(A)$ is a quasi-isomorphism when $A$ is a Koszul algebra, we get the following result:
Proposition 5.5.
Let $A$ be a Koszul algebra and $M$ be a graded $A$-bimodule. The canonical inclusion of chain complexes $K(A)\hookrightarrow B(A)$ induces an isomorphism between the graded Hochschild cohomology of $A$ with coefficients in $M$ and the cohomology of the cochain complex $(\underline{\mathrm{Hom}}_{A^{e}}(K(A),M),d^{*})$ where $d^{*}$ denotes the differential induced by the differential $d$ of the chcain complex $K(A)$. ∎
We now show that for a Koszul algebra an even simpler complex can be used to compute its Hochschild cohomology.
Definition 5.6.
Let $A$ be a Koszul algebra and $M$ be a graded $A$-bimodule. For every $n\geq 0$ and $s\in\mathbb{Z}$, let $\partial$ denote the $\mathbb{F}_{2}$-linear map
$\partial_{n}\colon\mathrm{Hom}_{\mathbb{F}_{2}}(K_{n}^{n}(A),M_{n+s})\to\mathrm{Hom}_{\mathbb{F}_{2}}(K_{n+1}^{n+1}(A),M_{n+1+s})$
defined by setting
$\partial_{n}(f)(x_{1}\otimes\ldots\otimes x_{n+1})=x_{1}f(x_{2}\otimes\ldots\otimes x_{n+1})+f(x_{1}\otimes\ldots\otimes x_{n})x_{n+1}.$
It is easily verified that $\partial\circ\partial=0$. We refer to the cochain complex
(13) $(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K^{*}_{*}(A),M[s]),\partial)$
as the simplified Koszul cochain complex.
Let $A$ be a Koszul algebra and $M$ be a graded $A$-bimodule. For every $n$ and every integer $s$, isomorphism (11) descends to an isomorphism
(14) $\underline{\mathrm{Hom}}_{A^{e}}(K_{n}(A),M[s])\xrightarrow{\cong}\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K_{n}^{n}(A),M_{n+s}).$
It is straight-forward to check that isomorphism (14) yields an isomorphism of cochain complexes
(15) $(\underline{\mathrm{Hom}}_{A^{e}}(K_{*}(A),M[s]),d^{*})\xrightarrow{\cong}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K_{*}^{*}(A),M[s]),\partial).$
By Proposition 5.5, the inclusion $K(A)\hookrightarrow B(A)$ induces an isomorphism between the Hochschild cohomology group $\mathrm{HH}^{n,s}(A,M)$ and the $n$th-cohomology of the left-hand side of (15). Since (15) is an isomorphism of cochain complexes, this proves the following result:
Proposition 5.7.
Let $A$ be a Koszul algebra and $M$ be a graded $A$-bimodule. For every pair $(n,s)$, the composition of the restriction along the inclusion $K(A)\hookrightarrow B(A)$ and isomorphism (15) defines an isomorphism
(16) $\mathrm{HH}^{n,s}(A,M)\xrightarrow{\cong}H^{n}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K_{*}^{*}(A),M[s]),\partial)$
from the graded Hochschild cohomology group $\mathrm{HH}^{n,s}(A,M)$ to the $n$th cohomology group of the cochain complex $(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K^{*}_{*}(A),M[s]),\partial)$. ∎
Theorem 4.8 and Proposition 5.7. then imply the following criterion for intrinsic formality:
Corollary 5.8.
Let $A$ be a Koszul algebra over $\mathbb{F}_{2}$. Assume that
$H^{n}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K_{*}^{*}(A),A[2-n]),\partial)=0~{}\text{for all}~{}n\geq 3.$
Then $A$ is intrinsically formal. ∎
Finally, to prove our main technical result in the next sections we will apply the following spectral sequence. We will provide a proof of the following proposition in appendix B.
Proposition 5.9.
[…]
6. Boolean and dual graded algebras
In this section we introduce the two types of graded algebras for which we will compute the Hochschild cohomology of their connected sums in the next sections. For more details and background on Boolean rings we refer to appendix A.
Definition 6.1.
We say that a ring $B$ is Boolean if $x^{2}=x$ for every $x\in B$. By Stone’s representation theorem every such $B$ is a subring of $\mathbb{F}_{2}^{X}$ with $X$ denoting the set of maximal ideals, see also Proposition A.3. For every Boolean ring $B$, let $B_{*}=\bigoplus_{n\geq 0}B_{n}$ denote the graded ring such that $B_{0}=\mathbb{F}_{2}$ and for every $n\geq 1$ we set $B_{n}=B$. For every pair of positive integers $m,n$ the multiplication $B_{m}\times B_{n}\to B_{m+n}$ is the multiplication in the ring $B=B_{m}=B_{n}=B_{m+n}$. We call $B_{*}$ the associated Boolean graded algebra.
For the next lemma, we recall that the polynomial ring $A=\mathbb{F}_{2}[x]$ is a quadratic algebra with $A_{1}=\mathbb{F}_{2}\langle x\rangle$, the one-dimensional $\mathbb{F}_{2}$-vector space generated by $x$, and trivial relations. For the next lemma, we recall from Corollary A.16 that every finite Boolean ring has an orthonormal basis.
Lemma 6.2.
Let $B$ be a finite Boolean ring. Let $x_{1},x_{2},\ldots,x_{n}$ be the orthonormal basis of $B$, i.e., $x_{1},x_{2},\ldots,x_{n}$ form a basis of $B$ as a vector space over $\mathbb{F}_{2}$ and $x_{i}x_{j}=0$ for $i\neq j$. Then the associated Boolean graded algebra $B_{*}$ is a quadratic algebra and is isomorphic to the connected sum of the $\mathbb{F}_{2}[x_{j}]$, i.e.,
$B_{*}\cong\mathbb{F}_{2}[x_{1}]\sqcap\cdots\sqcap\mathbb{F}_{2}[x_{n}].$
Moreover, $B_{*}$ is a Koszul algebra.
Proof.
[…]
is exact, and $\mu\colon K(A)\to A$ is a quasi-isomorphism of chain complexes. Thus, $A$ is Koszul and hence so is $B_{*}$. ∎
Proposition 6.3.
Let $B$ be a Boolean ring and let $B_{*}$ be its assoctiated Boolean graded algebra. Then $B_{*}$ is a Koszul algebra.
Proof.
This follows from the fact that every Boolean ring is the inductive limit of finite Boolean rings. Indeed, every ring is the inductive limit of its finitely generated subrings. Every subring of a Boolean ring is Boolean, and from Stone’s representation theorem it is also clear that every finitely generated Boolean ring is finite. Also note that the functor $B\mapsto B_{*}$ of Definition 6.1 maps injective morphisms to injective morphisms, so for every Boolean ring $B$ the graded $\mathbb{F}_{2}$-algebra $B_{*}$ is the inductive limit of graded $\mathbb{F}_{2}$-algebras of the form $A_{*}$ where $A\subseteq B$ is a finite Boolean ring. By Lemma 6.2, it now suffices to observe that the property of being a Koszul algebra is preserved under passing to inductive limits in the category of graded algebras over a field. This was shown by Bruns and Gubeladze in [6, Lemma 1.1, page 48]. ∎
As a generalisation of the algebra of dual numbers $k[\varepsilon]/(\varepsilon^{2})$, we refer to the following type of graded algebras as dual algebras (even though we do not consider them as the dual of another algebra).
Definition 6.4.
Let $V$ be a vector space over $\mathbb{F}_{2}$. We call the non-negatively graded $\mathbb{F}_{2}$-algebra $V_{*}=\bigoplus_{j\geq 0}V_{j}$ with $V_{0}=\mathbb{F}_{2}$, $V_{1}=V$ and $V_{j}=0$ for $j\geq 2$ the associated dual algebra. A dual algebra $V_{*}$ is quadratic with relations $R=V_{1}\otimes_{\mathbb{F}_{2}}V_{1}$.
Lemma 6.5.
For every $\mathbb{F}_{2}$-vector space $V$, the dual algebra $V_{*}$ is Koszul. For $k+m\geq 2$, we have $\mathrm{HH}^{k,m}(V_{*})=0$.
Proof.
That $V_{*}$ is Koszul follows from the fact that the inclusion map of $K(V_{*})$ into the bar complex is the identity. The vanishing of $\mathrm{HH}^{k,m}(V_{*})$ for $k+m\geq 2$ follows from the fact that there are no nontrivial cochains in the Koszul cochain complex for $k+m\geq 2$, since $V_{j}=0$ for $j\geq 2$. ∎
Proposition 6.6.
Let $V$ be an $\mathbb{F}_{2}$-vector space, let $V_{*}$ be its associated dual algebra, $B$ be a Boolean ring and let $B_{*}$ be its associated Boolean graded algebra. Then $V_{*}\sqcap B_{*}$ is a Koszul algebra.
Proof.
By Lemma 6.5 we know that $V_{*}$ is Koszul, and by Proposition 6.3 we know that $B_{*}$ is Koszul. By [27, Chapter 3, Corollary 1.2 on page 58] the connected sum of Koszul algebras is Koszul. This proves the assertion. ∎
7. Kadeishvili’s vanishing for connected sums of dual and Boolean graded algebras - the finite case
In this section we are going to prove that connected sums of dual algebras and the Boolean graded algebras associated to a Boolean ring have vanishing Hochschild cohomology when we restrict to a finite subring. In the next section we extend the computation to the case of arbitrary Boolean subrings.
We fix an $\mathbb{F}_{2}$-vector space $V$, and let $V_{*}$ be its associated dual algebra. We also fix a Boolean ring $B$ and a finite Boolean subring $A\subset B$. Let $A_{*}$ and $B_{*}$ be the associated Boolean graded algebras. Our next goal is to compute the groups $\mathrm{HH}^{k,m}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})$ for $m<0$.
By Propositions 5.7 and 6.6, we can do this by using the simplified Koszul cochain complex (13) of $V_{*}\sqcap A_{*}$ of Definition 5.6, i.e., we consider the cochain complex
$(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K^{*}_{*}(V_{*}\sqcap A_{*}),(V_{*}\sqcap B_{*})[m]),\partial).$
First we determine the $\mathbb{F}_{2}$-vector spaces $K^{k}_{k}(V_{*}\sqcap A_{*})$. This will require some preparation.
Definition 7.1.
Let $I$ denote a basis of $V$. Let $x_{1},x_{2},\ldots,x_{n}$ be the orthonormal basis of $A$, see Corollary A.16. Let $J$ denote the set $\{x_{1},x_{2},\ldots,x_{n}\}$. A sequence $t_{1},t_{2},\ldots,t_{k}\in I\cup J$ is admissible if for every $j=1,2,\ldots,k-1$ such that $t_{j},t_{j+1}\in J$ we have $t_{j}\neq t_{j+1}$. We say that an admissible sequence $t_{1},t_{2},\ldots,t_{k}\in I\cup J$ is stable if $t_{1},t_{k}\in J$ and $t_{1}=t_{k}$, and it is unstable otherwise.
Notation 7.2.
For every sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$, let the same symbol $\mathbf{t}$ denote the element $t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k}\in K^{k}_{k}(V_{*}\sqcap A_{*})$ by slight abuse of notation.
Lemma 7.3.
Write $Q_{*}=V_{*}\sqcap A_{*}$. The subspace $K^{k}_{k}(Q)\subset Q_{1}^{\otimes k}$ is spanned by the basis consisting of the vectors $\mathbf{t}$, where $\mathbf{t}$ is any admissible sequence.
Proof.
Let $R_{Q}\subseteq Q_{1}\otimes Q_{1}$ denote the relations in the tensor algebra defining the quadratic algebra $Q$. Note that the vectors $\mathbf{t}$, where $\mathbf{t}$ is any sequence in $I\cup J$, are linearly independent, so any subset of them will form a basis of their span. Therefore it will be sufficient to prove that for every index $j=1,2,\ldots,k-1$ the subspace:
$Q_{1}^{\otimes{j-1}}\otimes R_{Q}\otimes Q_{1}^{\otimes{k-j-1}}\subset Q_{1}^{\otimes k}$
is spanned by the vectors $\mathbf{t}$, where $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$ satisfies the condition $t_{j}\neq t_{j+1}$ if $t_{j},t_{j+1}\in J$. Let $T_{Q}\subset Q_{1}^{\otimes 2}$ be the subspace spanned by the vectors $x_{j}\otimes x_{j}$, where $j=1,2,\ldots,n$. Then $Q_{1}^{\otimes 2}=R_{Q}\oplus T_{Q}$, and hence
(17) $Q_{1}^{\otimes k}=\left(Q_{1}^{\otimes{j-1}}\otimes R_{Q}\otimes Q_{1}^{\otimes{k-j-1}}\right)\oplus\left(Q_{1}^{\otimes{j-1}}\otimes T_{Q}\otimes Q_{1}^{\otimes{k-j-1}}\right).$
Now if
$\underline{t}=\sum_{\mathbf{t}}a_{\mathbf{t}}\mathbf{t}=\left(\sum_{\{t_{j},t_{j+1}\}\nsubseteq J}a_{\mathbf{t}}\mathbf{t}+\sum_{\begin{subarray}{c}t_{j},t_{j+1}\in J\\ t_{j}\neq t_{j+1}\end{subarray}}a_{\mathbf{t}}\mathbf{t}\right)+\sum_{\begin{subarray}{c}t_{j},t_{j+1}\in J\\ t_{j}=t_{j+1}\end{subarray}}a_{\mathbf{t}}\mathbf{t}\quad(a_{\mathbf{t}}\in\mathbb{F}_{2})$
lies in $Q_{1}^{\otimes{j-1}}\otimes R_{Q}\otimes Q_{1}^{\otimes{k-j-1}}$, then the sum outside of the parentheses must be zero, since it lies in the second summand of the decomposition in (17), while the sum within the parentheses lies in the first summand of the decomposition in (17). The claim is now clear. ∎
We are going to use the following fact about Boolean rings.
Lemma 7.4.
Let $x,y\in B$. Let $(x)$, $(y)$ and $(x,y)$ denote the ideals in $B$ generated by $x$, $y$, $x$ and $y$, respectively. Then the following assertions hold:
  1. $(i)$
    for every $z\in B$, we have $z\in(x,y)$ if and only if $(1+x+y+xy)z=0$;
  2. $(ii)$
    if $xy=0$ then $(x,y)$ is the direct sum of $(x)$ and $(y)$.
In particular, for every $x,z\in B$ we have $z\in(x)$ if and only if $(1+x)z=0$.
Proof.
If $(1+x+y+xy)z=0$ then $z=z(x+y+xy)\in(x,y)$. If $z\in(x,y)$ then there are $a,b\in B$ such that $z=ax+by$, so
$(1+x+y+xy)z=(ax+by)+(ax+bxy)+(axy+by)+(axy+bxy)=0,$
so claim $(i)$ is true. Now assume that $xy=0$. Note that $(x,y)=(x)+(y)$, so we only need to show that $(x)\cap(y)=0$. If $z\in(x)\cap(y)$ then there are $a,b\in B$ such that $z=ax=by$. Therefore
$z=z^{2}=axby=abxy=0,$
so claim $(ii)$ holds, too. ∎
Now we assume that $j\geq 1$ and that $f\colon K^{k}_{k}(V_{*}\sqcap A_{*})\to B_{j}$ is a cocycle in the simplified Koszul cochain complex $(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(K^{*}_{*}(V_{*}\sqcap A_{*}),B_{*}[j-k]),\partial)$ of Definition 5.6. For the sake of simple notation, we will identify $B_{j}$ with $B_{1}$ in all that follows. Let $p\colon V_{1}\oplus B_{1}\to B_{1}$ be the unique linear map which is zero on $V_{1}$ and the identity on $B_{1}$.
Lemma 7.5.
For every admissible sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$ we have:
$f(\mathbf{t})\in(p(t_{1}),p(t_{k})).$
Proof.
We define the element $x:=1+p(t_{1})+p(t_{k})+p(t_{1})p(t_{k})\in A_{1}\subset B_{1}=B$. We claim that $x$ satisfies
$xt_{1}=0=xt_{k}.$
First, elements in $I$ are orthogonal to all elements and the claim is true for $t_{1},t_{k}\in I$. Second, if $t_{1}\in J$, then $xt_{1}=t_{1}+t_{1}^{2}+t_{1}p(t_{k})+t_{1}^{2}p(t_{k})=t_{1}+t_{1}+t_{1}p(t_{k})+t_{1}p(t_{k})=0$. And we have a similar calculation if $t_{k}\in J$. This shows the claim.
Hence $x\otimes t_{1}\otimes\cdots\otimes t_{k-1}\otimes t_{k}\in K^{k+1}_{k+1}(V_{*}\sqcap A_{*})$. Since $f$ is a cocycle:
$xf(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k})+f(x\otimes t_{1}\otimes\cdots\otimes t_{k-1})t_{k}=0.$
Multiplying both sides of this equation by $x$ we get that
$0=x^{2}f(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k})+f(x\otimes t_{1}\otimes\cdots\otimes t_{k-1})xt_{k}=xf(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k}).$
The claim now follows from part $(i)$ of Lemma 7.4. ∎
Definition 7.6.
For every stable admissible sequence $\mathbf{t}$, by Lemma 7.5, we have
$f(\mathbf{t})=\alpha(\mathbf{t})$
for a unique $\alpha(\mathbf{t})\in(p(t_{1}))$. Set $\beta(\mathbf{t})=\alpha(\mathbf{t})$ in this case. For every unstable admissible sequence $\mathbf{t}$, Lemma 7.5 implies
$f(\mathbf{t})=\alpha(\mathbf{t})+\beta(\mathbf{t}),\quad\alpha(\mathbf{t})\in(p(t_{1})),\beta(\mathbf{t})\in(p(t_{k})),$
where $\alpha(\mathbf{t}),\beta(\mathbf{t})$ are uniquely determined by part $(ii)$ of Lemma 7.4. For every admissible sequence $\mathbf{t}$ we call $\alpha(\mathbf{t}),\beta(\mathbf{t})$ the head, tail value of $f(\mathbf{t})$, respectively.
Let $A_{k}(I,n)$ denote the set of admissible sequences $t_{1},t_{2},\ldots,t_{k}\in I\cup J$, where we recall that $n$ denotes the cardinality of $J$, i.e., the $\mathbb{F}_{2}$-dimension of the finite subring $A\subset B$, which determines the associated Boolean graded algebra $A_{*}$ by Lemma 6.2.
Definition 7.7.
We define two maps $L\colon A_{k}(I,n)\to A_{k}(I,n)$ and $R\colon A_{k}(I,n)\to A_{k}(I,n)$, the left and right translations, as follows. Let
$L(t_{1},t_{2},\ldots,t_{k})=\begin{cases}t_{1},t_{2},\ldots,t_{k}&\text{, if the sequence is stable},\\ t_{2},t_{3},\ldots,t_{k},t_{1}&\text{, if the sequence is unstable, }\end{cases}$
and similarly
$R(t_{1},t_{2},\ldots,t_{k})=\begin{cases}t_{1},t_{2},\ldots,t_{k}&\text{, if the sequence is stable},\\ t_{k},t_{1},t_{2},\ldots,t_{k-1}&\text{, if the sequence is unstable. }\end{cases}$
These two maps translate the sequence to the left or to the right, respectively, and put the term which drops out to the other end. It is also clear that they are well-defined, i.e., they map admissible sequences to admissible sequences. Note that $L$ and $R$ are inverses of each other, so we get a permutation, or a cyclic group action on $A_{k}(I,n)$. The fixed points of this action are exactly the stable admissible sequences.
Lemma 7.8.
For every admissible sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$ we have:
$\alpha(R(\mathbf{t}))=\beta(\mathbf{t}),$
and equivalently:
$\beta(L(\mathbf{t}))=\alpha(\mathbf{t}).$
Proof.
The case of a stable admissible sequence is trivial. If $\mathbf{t}$ is unstable, the two identities above follow from the two relations:
$t_{k}f(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k})+f(t_{k}\otimes t_{1}\otimes\cdots\otimes t_{k-1})t_{k}=0,$
$t_{1}f(t_{2}\otimes\cdots\otimes t_{k}\otimes t_{1})+f(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k})t_{1}=0,$
respectively. ∎
Remark 7.9.
Note that the lemma above covers all relations between head and tail values for cocycles. This is true because all relations for cocycles are spanned by the relations:
(18) $t_{1}f(t_{2}\otimes t_{3}\otimes\cdots\otimes t_{k+1})+f(t_{1}\otimes t_{2}\otimes\cdots\otimes t_{k})t_{k+1}=0,$
where $\mathbf{t}=t_{1},t_{2},\ldots,t_{k+1}\in I\cup J$ is any admissible sequence. In particular there is no relation between head and tail values for admissible sequences lying in different orbits of $R$. Within every orbit of the right translation operator the head value of the right translate is equal to the tail value of the translated sequence, but otherwise we are free to choose the head and tail values within an orbit.
Definition 7.10.
We define two maps
$l\colon A_{k}(I,n)\to A_{k-1}(I,n)\quad\textrm{and}\quad r\colon A_{k}(I,n)\to A_{k-1}(I,n),$
the left and right truncations, as follows. Let
$l(t_{1},t_{2},\ldots,t_{k})=t_{1},t_{2},\ldots,t_{k-1},$
and similarly
$r(t_{1},t_{2},\ldots,t_{k})=t_{2},t_{3},\ldots,t_{k}.$
These maps are clearly well-defined, i.e., they map admissible sequences to admissible sequences. We say that an orbit $O\subset A_{k}(I,n)$ of $R$ is stable if it consists of one element, and it is unstable if it consists of at least two elements.
Lemma 7.11.
Let $O\subset A_{k}(I,n)$ be an unstable orbit and let $\mathbf{i},\mathbf{j}$ be two elements of $O$. The following holds:
  1. $(i)$
    if $l(\mathbf{i})=l(\mathbf{j})$ then $\mathbf{i}=\mathbf{j}$,
  2. $(ii)$
    if $r(\mathbf{i})=r(\mathbf{j})$ then $\mathbf{i}=\mathbf{j}$,
  3. $(iii)$
    we have $r(\mathbf{i})=l(\mathbf{j})$ if and only if $\mathbf{i}=R(\mathbf{j})$ and $\mathbf{j}=L(\mathbf{i})$.
Proof.
We first prove part $(i)$. Since $\mathbf{j}$ is a permutation of $\mathbf{i}$, the number of occurrences of $i_{k}$ in $\mathbf{j}$ is the same as in $\mathbf{i}$. Since $l(\mathbf{i})=l(\mathbf{j})$, the number of occurrences of $i_{k}$ in $l(\mathbf{j})$ is the same as in $l(\mathbf{i})$. This is only possible if $j_{k}=i_{k}$, and so part $(i)$ is true. The proof of part $(ii)$ is essentially the same as the proof of part $(i)$ above, just left and right have to be exchanged. By definition $r(R(\mathbf{j}))=l(\mathbf{j})$ and $l(L(\mathbf{i}))=r(\mathbf{i})$, so the reverse implication of part $(iii)$ is clearly true. On the other hand if $r(\mathbf{i})=l(\mathbf{j})$ then $r(\mathbf{i})=r(R(\mathbf{j}))$ by the above, so $\mathbf{i}=R(\mathbf{j})$ by part $(ii)$, and hence $\mathbf{j}=L(\mathbf{i})$, too. So part $(iii)$ holds, too. ∎
Definition 7.12.
The truncation set $\mathbb{T}(O)$ of an unstable orbit $O$ as above is its image with respect to the map $l$. By part $(iii)$ of Lemma 7.11 this is the same as the image of $O$ with respect to the map $r$. By parts $(i)$ and $(ii)$ of Lemma 7.11, respectively, we get that both $l:O\to\mathbb{T}(O)$ and $r\colon O\to\mathbb{T}(O)$ are bijective. Let $l^{-1}\colon\mathbb{T}(O)\to O$ and $r^{-1}\colon\mathbb{T}(O)\to O$ be the inverse of these two maps, respectively.
We are finally ready to prove the following theorem:
Theorem 7.13.
Let $V$ be an $\mathbb{F}_{2}$-vector space and $V_{*}$ its associated dual algebra. Let $B$ be a Boolean ring, $A\subset B$ a finite Boolean subring, and let $A_{*}$ and $B_{*}$, respectively, be the associated Boolean graded algebras. We have:
$\mathrm{HH}^{k,2-k}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})=0~{}\text{for all}~{}k\geq 3.$
Assume $|A|\geq 8$ and let $m<0$ be a negative integer. Then we have:
$\mathrm{HH}^{k,m}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})=0~{}\text{for all}~{}k\geq 0~{}\text{with}~{}k\neq 1-m.$
Proof.
For $k\geq 2$ and $j\geq 2$, we are going to show that every cocycle
$f\colon K^{k}_{k}(V_{*}\sqcap A_{*})\to B_{j}$
is a coboundary. This will imply that $\mathrm{HH}^{k,s}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})=0$ for all pairs $(k,m)$ with $k\geq 2$ and $m\geq 2-k$. The remaining cases will be discussed afterwards. By the above it will be enough to show the claim orbit by orbit, that is, we may assume without loss of generality that $f$ is supported on an orbit $O\subset A_{k}(I,n)$ of the right translation $R$. We will let $\alpha$ and $\beta$ denote the head and tail values of $f$, as above.
First assume that $O$ is a stable orbit $\{\mathbf{t}\}$ with $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$. In particular, $t_{1}=t_{k}$ and these two elements are in $J$. Let $g\colon K^{k-1}_{k-1}(V_{*}\sqcap A_{*})\to B_{j-1}$ be the cochain given by the rule:
$g(\mathbf{s})=\begin{cases}\alpha(\mathbf{t})t_{1}^{j-1}&\text{, if $\mathbf{s}=r(\mathbf{t})$},\\ 0&\text{, otherwise.}\end{cases}$
Lemma 7.14.
We have $\partial g=f$.
Proof.
Note that $t_{k-1}\neq t_{k}$ since $t_{k}\in J$ and $\mathbf{t}$ is admissible. Therefore $l(\mathbf{t})=t_{1},\ldots,t_{k-1}$ is different from $r(\mathbf{t})$, and hence
$\partial g(\mathbf{t})=t_{1}g(r(\mathbf{t}))+g(l(\mathbf{t}))t_{k}=\alpha(\mathbf{t})t_{1}^{j}+0=f(\mathbf{t}).$
Now let $\mathbf{s}=s_{1},s_{2},\ldots,s_{k}\in I\cup J$ be an admissible sequence different from $\mathbf{t}$. If $g(r(\mathbf{s}))\neq 0$ then $r(\mathbf{s})=r(\mathbf{t})$, and hence $s_{1}\neq t_{1}$, as $\mathbf{s}$ and $\mathbf{t}$ are different. Therefore $s_{1}g(r(\mathbf{s}))=s_{1}\alpha(\mathbf{t})t_{1}^{j-1}=0$ in this case since $j-1\geq 1$ by assumption, $\alpha(\mathbf{t})\in(p(t_{1}))$ and the product of $s_{1}$ and $t_{1}$ is zero when $s_{1}\neq t_{1}$. Otherwise $g(r(\mathbf{s}))=0$, and hence $s_{1}g(r(\mathbf{s}))=0$ trivially in this case.
If $g(l(\mathbf{s}))\neq 0$ then $l(\mathbf{s})=r(\mathbf{t})$, and so $s_{k-1}=t_{k}=t_{1}$. So $s_{k}$ is different from $t_{1}$, since $\mathbf{s}$ is admissible. Hence we have $g(l(\mathbf{s}))s_{k}=\alpha(\mathbf{t})t_{1}^{j-1}s_{k}=0$ in this case since $j-1\geq 1$ and the product of $t_{1}$ and $s_{k}$ is zero when $s_{k}\neq t_{1}$. Otherwise $g(l(\mathbf{s}))=0$, and hence $g(l(\mathbf{s}))s_{k}=0$ trivially in this case. We get that
$\partial g(\mathbf{s})=s_{1}g(r(\mathbf{s}))+g(l(\mathbf{s}))s_{k}=0+0=0.\qed$
Now assume that $O$ is an unstable orbit. For every sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup J$, let $a(\mathbf{t}),b(\mathbf{t})$ denote the first term $t_{1}$ and the last term $t_{k}$ of $\mathbf{t}$, respectively. Let $g\colon K^{k-1}_{k-1}(V_{*}\sqcap A_{*})\to B_{j-1}$ be the cochain given by the rule:
$g(\mathbf{s})=\begin{cases}\alpha(r^{-1}(\mathbf{s}))p(a(r^{-1}(\mathbf{s})))^{j-1}&\text{, if $\mathbf{s}\in\mathbb{T}(O)$},\\ 0&\text{, otherwise.}\end{cases}$
In particular, $g$ is supported on the truncation set of $O$.
Lemma 7.15.
We have:
$g(\mathbf{s})=\beta(l^{-1}(\mathbf{s}))p(b(l^{-1}(\mathbf{s})))^{j-1}$
for every $\mathbf{s}\in\mathbb{T}(O)$.
Proof.
Since $l(l^{-1}(\mathbf{s}))=\mathbf{s}=r(r^{-1}(\mathbf{s}))$ we get that $l^{-1}(\mathbf{s})=L(r^{-1}(\mathbf{s}))$ by part $(iii)$ of Lemma 7.11. Therefore $b(l^{-1}(\mathbf{s}))=a(r^{-1}(\mathbf{s}))$ and $\beta(l^{-1}(\mathbf{s}))=\alpha(r^{-1}(\mathbf{s}))$ by Lemma 7.8. Therefore
$g(\mathbf{s})=\alpha(r^{-1}(\mathbf{s}))p(a(r^{-1}(\mathbf{s})))^{j-1}=\beta(l^{-1}(\mathbf{s}))p(b(l^{-1}(\mathbf{s})))^{j-1}.\qed$
Lemma 7.16.
We have $\partial g=f$.
Proof.
Let $\mathbf{s}=s_{1},s_{2},\ldots,s_{k}\in I\cup J$ be an admissible sequence. First assume that $\mathbf{s}\in O$. Then
$\partial g(\mathbf{s})=s_{1}g(r(\mathbf{s}))+g(l(\mathbf{s}))s_{k}=\alpha(\mathbf{s})p(s_{1})^{j}+\beta(\mathbf{s})p(s_{k})^{j}=f(\mathbf{s})$
where we use the definition of $g$ and Lemma 7.15 in rewriting the first and second summand, respectively. Now assume that $\mathbf{s}$ is not in $O$. If $g(r(\mathbf{s}))\neq 0$, then $r(\mathbf{s})=r(\mathbf{t})$ for some $\mathbf{t}\in O$. Then $s_{1}\neq t_{1}$, as $\mathbf{s}$ and $\mathbf{t}$ are different. Therefore $s_{1}g(r(\mathbf{s}))=s_{1}\alpha(\mathbf{t})p(t_{1})^{j-1}=0$ in this case since $j-1\geq 1$ by assumption and $s_{1}\neq t_{1}$. Otherwise $g(r(\mathbf{s}))=0$, and hence $s_{1}g(r(\mathbf{s}))=0$ trivially in this case.
If $g(l(\mathbf{s}))\neq 0$ then $l(\mathbf{s})=l(\mathbf{t})$ for some $\mathbf{t}\in O$. Then $s_{k}\neq t_{k}$, as $\mathbf{s}$ and $\mathbf{t}$ are different. Therefore, since $j-1\geq 1$ and $t_{k}\neq s_{k}$, we have $g(l(\mathbf{s}))s_{k}=\beta(\mathbf{t})p(t_{k})^{j-1}s_{k}=0$ by Lemma 7.15 in this case. Otherwise $g(l(\mathbf{s}))=0$, and hence $g(l(\mathbf{s}))s_{k}=0$ trivially in this case. We get that
$\partial g(\mathbf{s})=s_{1}g(r(\mathbf{s}))+g(l(\mathbf{s}))s_{k}=0+0=0.\qed$
This finishes the proof for $k\geq 2$ and $m\geq 2-k$. Since both summands and hence the direct sum $V_{j}\oplus B_{j}$ are trivial for negative $j$, there are no nontrivial cochains and hence no nontrivial cocycles for $k+m<0$. Hence it remains to check the values $k\geq 1$ and $m=-k$. For $m=-k$, the first nontrivial differentials in the cochain complex we use to compute Hochschild cohomology are
$\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap A_{*}),\mathbb{F}_{2})\xrightarrow{\partial_{k}}\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k+1}_{k+1}(V_{*}\sqcap A_{*}),V_{1}\oplus B_{1})\xrightarrow{\partial_{k+1}}\ldots$
Let $f\colon K^{k}_{k}(V_{*}\sqcap A_{*})\to\mathbb{F}_{2}$ be a cocycle. Let ${\mathbf{t}}=t_{i_{1}}\otimes\ldots\otimes t_{i_{k}}$ be an admissible sequence. Assuming that the set $I\cup J$ consists of at least three elements, there is a $t_{i_{k+1}}$ different from both $t_{i_{1}}$ and $t_{i_{k}}$. Then
$t_{i_{1}}f(t_{i_{2}}\otimes t_{i_{3}}\otimes\cdots\otimes t_{i_{k+1}})+f(t_{i_{1}}\otimes t_{i_{2}}\otimes\cdots\otimes t_{i_{k}})t_{i_{k+1}}=0$
is only possible if the coefficients of $t_{i_{1}}$ and $t_{i_{k+1}}$ in the above sum on the left-hand side, i.e., $f(t_{i_{2}}\otimes t_{i_{3}}\otimes\cdots\otimes t_{i_{k+1}})$ and $f(t_{i_{1}}\otimes t_{i_{2}}\otimes\cdots\otimes t_{i_{k}})t$, are both 00. In particular, we must have $f({\mathbf{t}})=0$. Thus there are no nontrivial cocycles in this case. This finishes the proof of Theorem 7.13. ∎
Remark 7.17.
We note that the assumption $|A|\geq 8$ cannot be dropped. For example, if $A_{*}=\mathbb{F}_{2}[x_{1}]\sqcap\mathbb{F}_{2}[x_{2}]$, then $\mathrm{HH}^{k,-k}(A_{*})\neq 0$ for even $k>0$.
8. Mittag-Leffler functors and transfinite recursion on subrings
In this section we recall and prove some results that we will need to extend the claim of Theorem 7.13 to infinite subrings in the next section. First, we recall some facts about Mittag-Leffler functors.
Definition 8.1.
Let $\mathcal{AB}$ denote the category of abelian groups. For every ordinal $\kappa$, by slight abuse of notation we let the same symbol denote the associated small category induced by an ordering. Let $F\colon\kappa^{op}\to\mathcal{AB}$ be a functor. We say that the functor $F$ is Mittag-Leffler if
  1. (ML1)
    for every $\alpha\in\beta\in\kappa$ the map $F_{\beta}\to F_{\alpha}$ is surjective,
  2. (ML2)
    for every limit ordinal $\alpha\in\kappa$ the map $F_{\alpha}\to\lim_{\beta\in\alpha}F_{\beta}$ is surjective.
Lemma 8.2.
If $F\colon\kappa^{op}\to\mathcal{AB}$ is a functor such that
  1. (ML0)
    for every $\alpha+1\in\kappa$ the map $F_{\alpha+1}\to F_{\alpha}$ is surjective,
  2. (ML2)
    for every limit ordinal $\alpha\in\kappa$ the map $F_{\alpha}\to\lim_{\beta\in\alpha}F_{\beta}$ is surjective.
Then $F$ is Mittag-Leffler.
Proof.
We only need to check condition (ML1) of Definition 8.1 for $F$. Note that we may assume without the loss of generality that $\alpha=\emptyset$ by replacing $\kappa$ with the unique ordinal isomorphic to the segment $[\alpha,\kappa)$ and $F$ with $F|_{[\alpha,\kappa)}$. For every $\alpha\leq\beta$ in $\kappa$, let $\phi_{\beta,\alpha}:F_{\beta}\to F_{\alpha}$ be the transition map, and for every limit ordinal $\beta\in\kappa$, let $\phi_{\beta}:F_{\beta}\to\lim_{\alpha\in\beta}F_{\alpha}$ be the limit of the transition maps $\phi_{\beta,\alpha}$. Now let $\beta\in\kappa$ and $x\in F_{\emptyset}$ be arbitrary. We are going to construct using transfinite recursion for every $\delta\leq\beta$ an element $x_{\delta}\in F_{\delta}$ with the following properties:
  1. $(i)$
    we have $x_{\emptyset}=x$,
  2. $(ii)$
    for every $\gamma\leq\delta\leq\beta$ we have $\phi_{\delta,\gamma}(x_{\delta})=x_{\gamma}$,
  3. $(iii)$
    for every limit ordinal $\delta\leq\beta$ we have $\phi_{\delta}(x_{\delta})=\lim_{\gamma\in\delta}x_{\gamma}$.
This is sufficient to conclude that claim, as $\phi_{\beta,\emptyset}(x_{\beta})=x_{\emptyset}=x$, so $\phi_{\beta,\emptyset}$ is surjective, as $x$ was arbitrary. Also note that the limit $\lim_{\gamma\in\delta}x_{\gamma}$ in $(iii)$ makes sense because of the compatibility in $(ii)$. Let us start the transfinite recursion: for the initial $\delta=\emptyset$ just take $x_{\emptyset}=x$. This satisfies $(i)$, the only non-trivial condition in this case. When $\delta$ is a limit ordinal there is an $x_{\delta}\in F_{\delta}$ such that $\phi_{\delta}(x_{\delta})=\lim_{\gamma\in\delta}x_{\gamma}$ by condition (ML2). Clearly $\phi_{\delta,\gamma}(x_{\delta})=x_{\gamma}$ for every $\gamma\in\delta$. When $\delta=\gamma+1$ is a successor ordinal then there is an $x_{\delta}\in F_{\delta}$ such that $\phi_{\delta,\gamma}(x_{\delta})=x_{\gamma}$ by condition (ML0). If $\alpha\in\delta$, then either $\alpha=\gamma$, and hence $\phi_{\delta,\alpha}(x_{\delta})=x_{\alpha}$ by the above, or $\alpha\in\gamma$, and hence $\phi_{\delta,\alpha}(x_{\delta})=\phi_{\gamma,\alpha}(\phi_{\delta,\gamma}(x_{\delta}))=\phi_{\gamma,\alpha}(x_{\gamma})=x_{\alpha}$ by the induction hypothesis. ∎
Proposition 8.3.
If $F\colon\kappa^{op}\to\mathcal{AB}$ is a Mittag-Leffler functor then $\underleftarrow{\lim}^{n}F=0$ for every $n>0$.
Proof.
This is Corollary A.3.14 of [25]. ∎
Lemma 8.4.
Let $F,G,H\colon\kappa^{op}\to\mathcal{AB}$ be functors which form a short exact sequence:
[…]
If $F$ and $G$ are Mittag–Leffler then $\underleftarrow{\lim}^{n}H=0$ for every $n>0$.
Proof.
Since the higher limit functors $\underleftarrow{\lim}^{n}$ are derived functors, we have a long cohomological exact sequence (see [39, §3.5]):
[…]
associated to the short exact sequence above. Now we only need to apply Proposition 8.3. ∎
Next we prove a result on the structure of the subrings of a Boolean ring. Let $B$ be a Boolean ring. For every subring $A\subset B$ and for every $x\in B$, let $A\langle x\rangle$ denote the subring of $B$ generated by $A$ and $x$. We assume now that $B$ is infinite and let $A\subset B$ be an infinite subring. Let $\kappa=|A|$ be its cardinality and choose a bijection $x\colon\kappa\to A$. Fix a subring $A_{0}\subset A$ such that $|A_{0}|=8$. Using transfinite recursion we can construct a set of subrings
$\{A_{\alpha}\subseteq A\mid\alpha\in\kappa\}$
of $A$ such that
  1. $(a)$
    we have $A_{\emptyset}=A_{0}$,
  2. $(b)$
    we have $A_{\alpha+1}=A_{\alpha}\langle x(\alpha)\rangle$ for every $\alpha\in\kappa$,
  3. $(c)$
    we have $\bigcup_{\beta\in\alpha}A_{\beta}=A_{\alpha}$ for every limit ordinal $\alpha\in\kappa$.
For this collection of subrings we have some additional properties:
Lemma 8.5.
The following holds:
  1. $(i)$
    we have $\bigcup_{\alpha\in\kappa}A_{\alpha}=A$,
  2. $(ii)$
    we have $A_{\beta}\subseteq A_{\alpha}$ for every $\beta\in\alpha\in\kappa$,
  3. $(iii)$
    we have $|A_{\alpha}|<\omega$ for every finite $\alpha\in\kappa$,
  4. $(iv)$
    we have $|A_{\alpha}|\leq|\alpha|$ for every infinite $\alpha\in\kappa$,
  5. $(v)$
    we have $8\leq|A_{\alpha}|<|A|$ for every $\alpha\in\kappa$.
Proof.
For every $\beta\in\kappa$ we have $x(\beta)\in A_{\beta}\langle x(\beta)\rangle=A_{\beta+1}\subseteq\bigcup_{\alpha\in\kappa}A_{\alpha}$, so the latter is $A$, and hence $(i)$ holds. We are going to show $(ii)$ by transfinite induction. When $\alpha$ is a limit ordinal then $A_{\beta}\subseteq\bigcup_{\beta\in\alpha}A_{\beta}=A_{\alpha}$ by condition $(c)$. When $\alpha=\gamma+1$ is a successor ordinal then $A_{\gamma}\subseteq A_{\gamma}\langle x_{\gamma}\rangle=A_{\alpha}$ by $(b)$. If $\beta\in\alpha$, then either $\beta=\gamma$, and hence $A_{\beta}\subseteq A_{\alpha}$ by the above, or $\beta\in\gamma$, and hence $A_{\beta}\subseteq A_{\gamma}\subseteq A_{\alpha}$ by the induction hypothesis. So $(ii)$ is true. Since every finitely generated subring of $A$ is finite, we get $(iii)$ by induction, as every positive finite ordinal is a successor ordinal, so condition $(b)$ applies.
Now we are going to show $(iv)$ by transfinite induction. When $\alpha=\omega$ or a limit ordinal more generally, we have $|A_{\beta}|\leq|\alpha|$ for every $\beta\in\alpha$ either by part $(iii)$ (when $\alpha=\omega$) or by the induction hypothesis. Therefore
$|A_{\alpha}|=|\bigcup_{\beta\in\alpha}A_{\beta}|\leq|\alpha|^{2}=|\alpha|$
using condition $(c)$. When $\alpha=\beta+1$ is a successor ordinal then by condition $(b)$ we have $A_{\alpha}=A_{\beta}\langle x_{\beta}\rangle$, so the latter is generated by $|\alpha|+1=|\alpha|$ elements, and hence its cardinality is at most $|\alpha|$. Therefore $(iv)$ holds, too. For every $\alpha\in\kappa$ we have $A_{\emptyset}\subseteq A_{\alpha}$ by part $(ii)$, so $8=|A_{\emptyset}|\leq|A_{\alpha}|$. On the other hand, $|A_{\alpha}|<|A|$ for every $\alpha\in\kappa$ by part $(iii)$, when $\alpha$ is finite, and by part $(iv)$, when $\alpha$ is infinite since $|A|=\kappa>\alpha$. Therefore $(v)$ is true. ∎
9. Kadeishvili’s vanishing for connected sums of dual and Boolean graded algebras - the infinite case
Our next goal is to extend Theorem 7.13 to the case of infinite subrings $A\subseteq B$. In particular, we will extent the theorem to the case $A=B$. This will require a considerable amount of work.
Notation 9.1.
We fix a vector space $V$ over $\mathbb{F}_{2}$ and let $V_{*}$ denote the associated dual algebra over $\mathbb{F}_{2}$. We fix a Boolean ring $B$ and let $B_{*}$ denote the associated graded Boolean algebra. For every subring $A\subseteq B$ with associated Boolean graded algebra $A_{*}$ consider the cochain complex:
[…]
Let $F^{k,d-k}(A)$ denote the group $\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k+d}_{k+d}(V_{*}\sqcap A_{*}),(V_{*}\sqcap B_{*})_{d})$ of degree $d$ cochains in this complex, and let $G^{k,d-k}(A)\subseteq F^{k,d-k}(A)$ be the subgroup of cocycles.
Lemma 9.2.
Let $A\subset B$ be a subring such that $|A|\geq 8$. Then $G^{k,-k}(A)=0$ for all $k\geq 1$.
Proof.
Since $|A|\geq 8$ there is a finite subring $A_{0}\subseteq A$ such that $|A_{0}|=8$. Since every finitely generated subring of $A$ is finite, the ring $A$ is the union of its finite subrings containing $A_{0}$. We already know the claim for the latter by the proof of Theorem 7.13. So by taking the limit we get the claim for $A$, too. ∎
As a consequence we get the following
Corollary 9.3.
Let $A\subset B$ be a subring such that $|A|\geq 8$. Then
$\mathrm{HH}^{k,-k+m}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})=0$
for all $k\geq 1,m\leq 0$.
Proof.
For $m=0$, the assertion follows from Lemma 9.2. For $m<0$, there are no nontrivial maps in $\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k+1}_{k+1}(V_{*}\sqcap A_{*}),V_{m}\oplus B_{m})$, since both $V_{m}$ and $B_{m}$ are trivial for negative $m$. ∎
For every subring $A\subset B$ and for every $x\in B$, let $A\langle x\rangle$ denote the subring of $B$ generated by $A$ and $x$.
Proposition 9.4.
Let $A\subset B$ be a subring. Then for every $x\in B$ the induced map $G^{k,1-k}(A\langle x\rangle)\to G^{k,1-k}(A)$ is surjective for all $k\geq 2$.
Proof.
Let $f\in G^{k,1-k}(A)$ be an arbitrary cocycle. We need to show that $f$ has a lift to $G^{k,1-k}(A\langle x\rangle)$. Note that for every subring $C\subset B$ every function in
$\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap C_{*}),V_{1})\subseteq\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap C_{*}),V_{1}\oplus B_{1})$
is a cocycle, since the multiplication with elements of degree $\geq 1$ in $V_{*}$ vanishes. Moreover, every $\mathbb{F}_{2}$-vector space homomorphism $K^{k}_{k}(V_{*}\sqcap A_{*})\to V_{1}$ can be factored through $K^{k}_{k}(V_{*}\sqcap A_{*})\to K^{k}_{k}(V_{*}\sqcap A\langle x\rangle_{*})$ by sending $x$ to zero. This implies that the map
$\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap A\langle x\rangle_{*}),V_{1})\to\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap A_{*}),V_{1})$
is surjective. Since every $\mathbb{F}_{2}$-vector space homomorphism $K^{k}_{k}(V_{*}\sqcap A_{*})\to V_{1}\oplus B_{1}$ can be written as a sum of homomorphisms into $V_{1}$ and $B_{1}$, respectively, it remains to show that cocycles with values in $B_{1}$ can be lifted.
We therefore assume from now on that the cocycle $f\in G^{k,1-k}(A)$ takes values in $B_{1}$. Consider $B_{1}$ as a $B$-module and write $f=xf+(1-x)f$. Note that $xf$ is again a cocycle, since multiplication of $x$ with elements in $V_{*}\sqcap A_{*}$ vanishes. The definition of the differential then yields that $xf$ is a cocycle. That implies that $(1-x)f$ is a cocycle as well. Therefore it will be enough to show that $xf$ and $(1-x)f$ have lifts to $G^{k,1-k}(A\langle x\rangle)$. Since the roles of $x$ and $1-x$ are symmetrical, we may assume without the loss of generality that $f=xf$. We are going to construct the lift when $A$ is finite first. We will need some preparations to do so. For more background on Boolean rings and proofs of the assertion on Boolean rings we refer to appendix A.
Notation 9.5.
For every Boolean ring $C$, let $\mathbf{S}(C)=\mathrm{Spec}(C)$ denote the spectrum of $C$. For every set $X$, let $\mathbf{B}(X)$ denote the ring of $\mathbb{F}_{2}$-valued functions on $X$. Note that every prime ideal in a Boolean ring $C$ is the kernel of a ring homomorphism $C\to\mathbb{F}_{2}$, see Proposition A.3 and Theorem A.10. So every element $x\in C$ gives rise to a function in $\mathbf{B}(\mathbf{S}(C))$ which furnishes a ring homomorphism $\mathfrak{f}\colon C\to\mathbf{B}(\mathbf{S}(C))$. This map is an isomorphism by Theorem A.6.
Remark 9.6.
We say that an element $x\in C$ is an atom if the support of $\mathfrak{f}(x)$ is a one element set, i.e., we have $\mathfrak{f}(x)(p)\neq 0$ for exactly one $p\in\mathbf{S}(C)$. From this perspective the unique orthonormal basis of a finite Boolean ring $C$ is just the collection of all atoms of $C$. Therefore we have a natural bijection between the elements of the unique orthonormal basis of $C$ and the elements of $\mathbf{S}(C)$, see Corollary A.16. We will identify these two sets in all that follows.
Definition 9.7.
For every homomorphism $h\colon C\to D$ between Boolean rings let $h^{*}\colon\mathbf{S}(D)\to\mathbf{S}(C)$ be the induced map. Let $a\colon A\to A\langle x\rangle$ be the inclusion map. Since $a$ is injective, the induced map $a^{*}\colon\mathbf{S}(A\langle x\rangle)\to\mathbf{S}(A)$ is surjective. Since every ring homomorphism $A\langle x\rangle\to\mathbb{F}_{2}$ is uniquely determined by its restriction to $A$ and $x$, the fibres of $a^{*}$ have at most two elements. Let $s\colon\mathbf{S}(A)\to\mathbf{S}(A\langle x\rangle)$ be the section of $a^{*}$ such that $s(p)$ is the unique element of $(a^{*})^{-1}(p)$, if $(a^{*})^{-1}(p)$ is a one element set, and $s(p)$ is the unique element of $(a^{*})^{-1}(p)$ such that $\mathfrak{f}(x)(s(p))=1$, otherwise, for every $p\in\mathbf{S}(A)$.
Notation 9.8.
Fix a basis $I$ of $V$. For every homomorphism $h\colon C\to D$ between Boolean rings let by slight abuse of notation $h^{*}$ also denote the map $I\cup\mathbf{S}(D)\to I\cup\mathbf{S}(C)$ which is the identity on $I$ and the map induced by $h$ restricted to $\mathbf{S}(D)$. Similarly let $s$ also denote the map $I\cup\mathbf{S}(A)\to I\cup\mathbf{S}(A\langle x\rangle)$ which is the identity on $I$ and the section introduced above, when restricted to $\mathbf{S}(A)$. Moreover for every sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(D)$ let $h^{*}(\mathbf{t})$ denote the sequence $h^{*}(t_{1}),h^{*}(t_{2}),\ldots,h^{*}(t_{k})\in I\cup\mathbf{S}(C)$, and for every sequence $t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A)$ let $s(\mathbf{t})$ denote the sequence $s(t_{1}),s(t_{2}),\ldots,s(t_{k})\in I\cup\mathbf{S}(A\langle x\rangle)$.
Definition 9.9.
Let $C$ be a subring of $B$. Under the identification in Remark 9.6 above, admissible sequences for $C$ are taking values in $I\cup\mathbf{S}(C)$. A sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A\langle x\rangle)$ is $x$-admissible if it is admissible and $\mathbf{t}=s(a^{*}(\mathbf{t}))$.
Lemma 9.10.
The following holds:
  1. $(i)$
    if $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A\langle x\rangle)$ is $x$-admissible then $a^{*}(\mathbf{t})$ is admissible,
  2. $(ii)$
    the maps $a^{*}$ and $s$ induce a bijection between the set of $x$-admissible sequences in $I\cup\mathbf{S}(A\langle x\rangle)$ and admissible sequences in $I\cup\mathbf{S}(A)$.
Proof.
If $a^{*}(\mathbf{t})$ were not admissible, there would be an index $i$ such that $a^{*}(t_{i})=a^{*}(t_{i+1})$. Since for every $u\in I\cup\mathbf{S}(A)$ the element $s(u)$ is unique, this means that $t_{i}=t_{i+1}$ which contradicts that $\mathbf{t}$ is admissible. So claim $(i)$ is true. By construction $s$ maps admissible sequences in $I\cup\mathbf{S}(A)$ to $x$-admissible sequences in $I\cup\mathbf{S}(A\langle x\rangle)$. Therefore $a^{*}$ and $s$ are maps between these sets, and since they are inverses of each other, we get that claim $(ii)$ holds, too. ∎
Now we are going to define a function $f_{x}\in\mathrm{Hom}_{\mathbb{F}_{2}}(K^{k}_{k}(V_{*}\sqcap A\langle x\rangle_{*}),B_{1})$ as the unique function such that for every admissible sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A\langle x\rangle)$ we have:
$f_{x}(x_{\mathbf{t}})=\begin{cases}f(x_{a^{*}(\mathbf{t})})&\text{, if $\mathbf{t}$ is $x$-admissible,}\\ 0&\text{, otherwise. }\end{cases}$
For every admissible sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A)$ we compute:
$f_{x}(x_{\mathbf{t}})=\sum_{\mathbf{u}\in(a^{*})^{-1}(\mathbf{t})}f_{x}(x_{\mathbf{u}})=f_{x}(x_{s(\mathbf{t})})=f(x_{\mathbf{t}}),$
where we used the additivity of $f_{x}$ and part $(ii)$ of Lemma 9.10 in the first equation, and its definition in the second and third equations. Therefore the restriction of $f_{x}$ onto $K^{k}_{k}(V_{*}\sqcap A_{*})$ is $f$. Since the map $a^{*}$ commutes with translations, and the translates of $x$-admissible sequences are $x$-admissible, in order to show that $f_{x}$ is a cocycle it will be sufficient to show for every $x$-admissible sequence $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A\langle x\rangle)$ that
$f_{x}(x_{\mathbf{t}})\in(t_{1},t_{k}).$
Note that $xt\in(s(t))$ for every $t\in I\cup\mathbf{S}(A)$. Therefore
$f_{x}(x_{\mathbf{t}})=f(x_{a^{*}(\mathbf{t})})=xf(x_{a^{*}(\mathbf{t})})\in x(a^{*}(t_{1}),a^{*}(t_{k}))\subseteq(xa^{*}(t_{1}),xa^{*}(t_{k}))\subseteq(t_{1},t_{k})$
using that $s(a^{*}(\mathbf{t}))=\mathbf{t}$. This concludes the proof when $A$ is finite.
Since every subring of a Boolean ring is Boolean, and every Boolean ring is the union of the directed set of its finite subrings, we need to show that the lifts we have constructed for finite subrings are compatible in order to prove the claim in general. Compatibility means the following: let $A\subseteq\widetilde{A}$ be a pair of finite subrings of $B$. Let $\widetilde{f}\in G^{k,1-k}(\widetilde{A})$ be a cocycle such that $\widetilde{f}=x\widetilde{f}$. (The latter condition implies that $\widetilde{f}$ takes values in $B_{1}$.) Let $f\in G^{k,1-k}(A)$ be the restriction of $\widetilde{f}$ onto $K^{k}_{k}(V_{*}\sqcap A_{*})$. Then $f_{x}$ is the restriction of the similarly constructed lift $\widetilde{f}_{x}$ of $\widetilde{f}$ onto $K^{k}_{k}(V_{*}\sqcap A\langle x\rangle_{*})$.
Notation 9.11.
Let $\widetilde{a}\colon\widetilde{A}\to\widetilde{A}\langle x\rangle$ be the inclusion map. Let $\widetilde{s}\colon I\cup\mathbf{S}(\widetilde{A})\to I\cup\mathbf{S}(\widetilde{A}\langle x\rangle)$ be the section of $\widetilde{a}^{*}$ defined the same way as $s$ and let the same symbol denote its extension to a map from series in $I\cup\mathbf{S}(\widetilde{A})$ to series in $I\cup\mathbf{S}(\widetilde{A}\langle x\rangle)$ by our usual abuse of notation. Let $i:A\to\widetilde{A}$ and $j:A\langle x\rangle\to\widetilde{A}\langle x\rangle$ be the inclusion maps.
Lemma 9.12.
The following holds:
  1. $(i)$
    if $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(\widetilde{A}\langle x\rangle)$ is $x$-admissible then $j^{*}(\mathbf{t})$ is $x$-admissible, too,
  2. $(ii)$
    for every admissible $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A)$ the maps $\widetilde{a}^{*}$ and $\widetilde{s}$ induce a bijection between the set of $x$-admissible sequences in $(j^{*})^{-1}(s(\mathbf{t}))$ and admissible sequences in $(i^{*})^{-1}(\mathbf{t})$.
Proof.
By part $(i)$ of Lemma 9.10 we already know that $j^{*}(\mathbf{t})$ is admissible. So if $a^{*}(\mathbf{t})$ were not $x$-admissible, there would be an index $i$ such that $j^{*}(t_{i})\neq s(a^{*}(j^{*}(t_{i})))$. Then $t_{i}\neq\widetilde{s}(\widetilde{a}^{*}(t_{i}))$ either, since $j^{*}\circ\widetilde{s}\circ\widetilde{a}^{*}=s\circ a^{*}\circ j^{*}$, which is a contradiction. So claim $(i)$ holds. Note that $\widetilde{a}^{*}\circ i^{*}=j^{*}\circ a^{*}$ and $\widetilde{s}\circ j^{*}=i^{*}\circ s$. Therefore $\widetilde{a}^{*}$ and $\widetilde{s}$ are maps between these sets, and since they are inverses of each other, we get that claim $(ii)$ holds, too. ∎
Now we are ready to show that $f_{x}$ is the restriction of $\widetilde{f}_{x}$. It will be sufficient to check this on $x_{\mathbf{t}}$ where $\mathbf{t}=t_{1},t_{2},\ldots,t_{k}\in I\cup\mathbf{S}(A\langle x\rangle)$ is admissible. If $\mathbf{t}$ is not $x$-admissible, then there is no $x$-admissible sequence in $(j^{*})^{-1}(\mathbf{t})$ by Lemma 9.12, so
$\widetilde{f}_{x}(x_{\mathbf{t}})=\sum_{\mathbf{u}\in(j^{*})^{-1}(\mathbf{t})}\widetilde{f}_{x}(x_{\mathbf{u}})=0=f_{x}(x_{\mathbf{t}}),$
where we used the additivity of $\widetilde{f}_{x}$ in the first equation, and the definition of $\widetilde{f}_{x}$ and $f_{x}$ in the second and third equations, respectively. If $\mathbf{t}$ is $x$-admissible, then
$\widetilde{f}_{x}(x_{\mathbf{t}})$ $=\sum_{\mathbf{u}\in(j^{*})^{-1}(\mathbf{t})}\widetilde{f}_{x}(x_{\mathbf{u}})=\sum_{\begin{subarray}{c}\mathbf{u}\in(j^{*})^{-1}(\mathbf{t})\\ \mathbf{u}=\widetilde{s}(\widetilde{a}^{*}(\mathbf{u}))\end{subarray}}\widetilde{f}(x_{\widetilde{a}^{*}(\mathbf{u})})$
$=\sum_{\mathbf{u}\in(i^{*})^{-1}(a^{*}(\mathbf{t}))}\widetilde{f}(x_{\mathbf{u}})=\widetilde{f}(x_{a^{*}(\mathbf{t})})=f(x_{a^{*}(\mathbf{t})})=f_{x}(x_{\mathbf{t}}),$
where we used the additivity of $\widetilde{f}_{x}$ in the first equation, the definition of $\widetilde{f}_{x}$ in the second equation, part $(ii)$ of Lemma 9.12 in the third equation, the additivity of $\widetilde{f}$ in the fourth equation, the fact that $f$ is the restriction of $\widetilde{f}$ in the fifth equation, and the definition of $f_{x}$ in the last equation. This finishes the proof of Proposition 9.4. ∎
We can now use Proposition 9.4 to show the vanishing of Hochschild cohomology groups. Let $A\subset B$ be an infinite subring. Let $\kappa=|A|$ be its cardinality and choose a bijection $x\colon\kappa\to A$. Let $\mathcal{SA}$ denote the category of subrings of $A$ with respect to the inclusions as morphisms. By part $(ii)$ of Lemma 8.5 we have a functor $\iota\colon\kappa\to\mathcal{SA}$ with $\iota(\alpha)=A_{\alpha}$ for every $\alpha\in\kappa$. Let $F^{k},G^{k},H^{k}\colon\kappa^{op}\to\mathcal{AB}$ be the composition of $\iota$ with the contravariant functors:
$C\mapsto F^{k-1,1-k}(C),\quad C\mapsto G^{k,1-k}(C),\quad C\mapsto H^{k}(C)=\mathrm{HH}^{k,1-k}(V_{*}\sqcap C_{*},V_{*}\sqcap B_{*})$
from $\mathcal{SA}$ to $\mathcal{AB}$, respectively.
Proposition 9.13.
We have $\underleftarrow{\lim}^{n}H^{k}=0$ for every $n>0$.
Proof.
By Lemma 9.2 there is a short exact sequence:
[…]
of functors. Therefore it will be enough show that both $F^{k}$ and $G^{k}$ are Mittag-Leffler by Lemma 8.4. Since $F^{k}$ is just the dual of an inductive system of vector spaces, it is Mittag-Leffler. As we can check that an element of $F^{k,1-k}(C)$ is a cocycle on $(k+2)$-uples of elements of $C$, we have $G^{k,1-k}(A_{\alpha})=\lim_{\beta\in\alpha}G^{k,1-k}(A_{\beta})$ for every limit ordinal $\alpha\in\kappa$, and hence $G^{k}$ satisfies property (ML2) of Definition 8.1 . By Proposition 9.4 the functor $G^{k}$ satisfies property (ML0) of Lemma 8.2. Hence $G^{k}$ is Mittag-Leffler by Lemma 8.2. ∎
Now we are ready to prove our main
Theorem 9.14.
Let $B$ be an infinite Boolean ring. Let $A\subseteq B$ be a subring such that $|A|\geq 8$. Let $m<0$ be a negative integer. Then
$\mathrm{HH}^{k,m}(V_{*}\sqcap A_{*},V_{*}\sqcap B_{*})=0$
for every $k\geq 0$ such that $k\neq 1-m$. In particular, we have
$\mathrm{HH}^{n,2-n}(V_{*}\sqcap B_{*})=0~{}\text{for all}~{}n\geq 3.$
Hence the graded algebra $V_{*}\sqcap B_{*}$ is intrinsically formal.
Proof.
We already know that this is true when $A$ is finite by Theorem 7.13, so we may assume without the loss of generality that $A$ is infinite. Let $\kappa=|A|$ be its cardinality. We prove Theorem 9.14 by transfinite induction on the cardinality $\kappa$ of $A$. This means we may assume that the claim holds for every subring of $A$ of cardinality less than $\kappa$. In particular the claim of the theorem holds for $A_{\alpha}$ for every $\alpha\in\kappa$ by part $(v)$ of Lemma 8.5. Fix an integer $m<0$. Since the higher limits of identically zero functors vanish, all rows of the spectral sequence of Proposition 5.9:
[…]
10. Proof of the main result
Let $G$ be a profinite group which is real projective. Let $\mathcal{X}(G)$ denote the conjugacy classes of involutions of $G$ equipped with its natural topology, and let $B=C(\mathcal{X}(\Gamma),\mathbb{F}_{2})$ be the ring of continuous functions from $\mathcal{X}(G)$ to $\mathbb{F}_{2}$ where we equip the latter with the discrete topology. Note that $B$ is a Boolean ring. Let $B_{*}$ be the associated Boolean graded algebra.
By Scheiderer’s Theorem 2.11, there is a surjective homomorphism of graded $\mathbb{F}_{2}$-algebras:
$\pi\colon H^{*}(G,\mathbb{F}_{2})\to B_{*}$
such that the $n$-th degree component:
$\pi_{n}\colon H^{n}(G,\mathbb{F}_{2})\to B_{n}$
is an isomorphism except perhaps when $n=1$, and it is surjective when $n=1$. Let $V\subseteq H^{1}(G,\mathbb{F}_{2})$ be the kernel of $\pi_{1}$, and let $V_{*}$ be the associated dual algebra over $\mathbb{F}_{2}$. By choosing a section
$\phi\colon B_{1}\to H^{1}(G,\mathbb{F}_{2})$
of $\pi_{1}$ we get an isomorphism of graded $\mathbb{F}_{2}$-algebras:
$H^{*}(G,\mathbb{F}_{2})\cong B_{*}\sqcap V_{*}.$
By Theorem 9.14, the connected sum $B_{*}\sqcap V_{*}$ satisfies Kadeishvili’s vanishing and is intrinsically formal. Thus $H^{*}(G,\mathbb{F}_{2})$ satisfies Kadeishvili’s vanishing and is intrinsically formal. By Corollary 4.9 this implies that the differential graded algebra $C^{*}(G,\mathbb{F}_{2})$ is formal. This finishes the proof of Theorem 1.3.
Finally, as explained previously, the Galois groups of fields with virtual cohomological dimension one are real projective profinite groups. Hence Scheiderer’s Theorem 2.11 on the structure of the cohomology algebra of real projective groups and Proposition 6.6 imply Theorem 1.11.
Appendix A Boolean rings
In this section we summarise the assertions on Boolean rings we use in Sections 6, 7 and 9.
Definition A.1.
A ring $R$ is called Boolean if $x^{2}=x$ for every $x\in R$.
Examples A.2.
The field with two elements $\mathbb{F}_{2}$ is a Boolean ring. In fact, since $x(1-x)=0$ for all $x$ in a Boolean ring, $\mathbb{F}_{2}$ is the only Boolean integral domain. The direct product of Boolean rings is Boolean, and so, for every set $X$, the direct product ring:
$\mathbb{F}_{2}^{X}=\prod_{i\in X}\mathbb{F}_{2}$
is a Boolean ring. Now let $X$ be a topological space, and let $\mathbf{B}(X)$ denote the ring of functions $f\colon X\to\mathbb{F}_{2}$ which are continuous with respect to the discrete topology on $\mathbb{F}_{2}$. Since the subrings of Boolean rings are Boolean, and $\mathbf{B}(X)$ is a subring of $\mathbb{F}_{2}^{X}$, we get that $\mathbf{B}(X)$ is Boolean, too.
Proposition A.3.
In a Boolean ring $R$ the following hold:
  1. $(i)$
    we have $02x=$ 0 for every $x\in R$,
  2. $(ii)$
    every prime ideal $\mathfrak{p}$ is maximal, and $R/\mathfrak{p}$ is the field with two elements,
  3. $(iii)$
    we have $(x,y)=(x+y-xy)$ for every $x,y\in R$,
  4. $(iv)$
    every finitely generated ideal is principal.
Proof.
Since
$2x=(2x)^{2}=4x^{2}=4x,$
we get that $02x=$ 0 by subtracting $2x$ from both sides. Now let $\mathfrak{p}$ be a prime ideal in $R$. Then the quotient $R/\mathfrak{p}$ is a Boolean ring. For every $x\in R/\mathfrak{p}$, we have $x(1-x)=0$ which implies that $x=0$ or $x=1$ since $R/\mathfrak{p}$ is an integral domain. Claim $(ii)$ follows. Note that
$x(x+y-xy)=x^{2}+xy-x^{2}y=x+xy-xy=x.$
Hence $x,y\in(x+y-xy)$. Since $x+y-xy\in(x,y)$, claim $(iii)$ is clear. Let $I=(x_{1},x_{2},\ldots,x_{n})$ be an finitely generated ideal of $R$. Since
$I=((x_{1},x_{2},\ldots,x_{n-1}),x_{n}),$
we may assume by induction on $n$ that $I=(x,y)$ for some $x,y\in R$. The claim now follows from part $(iii)$. ∎
Proposition A.4.
The spectrum $\mathrm{Spec}(R)$ of a Boolean ring is compact and totally separated.
Proof.
Since the spectrum of a commutative ring with a unity is compact, the same holds for $\mathrm{Spec}(R)$, too. Recall that a topological space $X$ is totally separated if for any two distinct points $x,y\in X$ there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$. Now let $\mathfrak{p},\mathfrak{q}\in\mathrm{Spec}(R)$ be two distinct points. Since they are maximal ideals, there is an $x\in R$ such that $x\in\mathfrak{p}$ and $x\not\in\mathfrak{q}$. Since $R/\mathfrak{p}=\mathbb{F}_{2}$ by part $(ii)$ of Proposition A.3, the former is equivalent to $1-x\not\in\mathfrak{p}$. As usual for every $f\in R$ let $D(f)\subseteq\mathrm{Spec}(R)$ denote the open subset
(19) $D(f)=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid f\not\in\mathfrak{p}\}.$
Then $\mathfrak{p}\in D(1-x),\mathfrak{q}\in D(x)$, the intersection $D(x)\cap D(1-x)$ is empty by part $(ii)$ of Proposition A.3, while the union $D(x)\cup D(1-x)$ is $\mathrm{Spec}(R)$, since if $x,1-x\in\mathfrak{m}$ for some $\mathfrak{m}\in\mathrm{Spec}(R)$ then $1\in\mathfrak{m}$ which is a contradiction. ∎
Notation A.5.
Let $R$ be a Boolean ring. Then for every $a\in R$ the corresponding section of the structure sheaf of $\mathrm{Spec}(R)$ is a continuous function
$\sigma(a)\colon\mathrm{Spec}(R)\to\mathbb{F}_{2}$
by part $(ii)$ of Proposition A.3, and the furnished map
$\sigma\colon R\to\mathbf{B}(\mathrm{Spec}(R))$
is a ring homomorphism.
Theorem A.6.
For every Boolean ring $R$, the map $\sigma\colon R\to\mathbf{B}(\mathrm{Spec}(R))$ is an isomorphism.
Proof.
For every $x\in R$ we have $x^{n}=x$ by induction, so if $x$ is nilpotent, then it is zero. Therefore the nilradical of $R$ is zero, so by Krull’s theorem $\sigma$ is injective. So we only need to show that $\sigma$ is surjective. Let $f\colon\mathrm{Spec}(R)\to\mathbb{F}_{2}$ be a continuous function. Then the set
$D(f)=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid f(x)=1\}$
is a closed subset, so it is compact. But it is also open, so it can be covered by open subsets of the form $D(a)$, where $a\in R$ by the definition of the topology of the spectrum of rings. Since $D(f)$ is compact, it can be covered by finitely many such, so
$D(f)=D(a_{1})\cup D(a_{2})\cup\cdots\cup D(a_{n})$
for some $a_{1},a_{2},\ldots,a_{n}\in R$. By part $(iv)$ of Proposition A.3, there is an $a\in R$ such that $(a)=(a_{1},\ldots,a_{n})$. Then
$D(a)$ =$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a\not\in\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a)\not\subseteq\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid(a_{1},\ldots,a_{n})\not\subseteq\mathfrak{p}\}$
=$=\{\mathfrak{p}\in\mathrm{Spec}(R)\mid a_{i}\not\in\mathfrak{p}\textrm{ for some $i$}\}$
$=D(a_{1})\cup D(a_{2})\cup\cdots\cup D(a_{n}),$
so $D(f)=D(a)$. Since $f$ and $a$ take values in $\mathbb{F}_{2}$, we get that $f=a$. ∎
Notation A.7.
Let $X$ be a topological space. Then for every $p\in X$ the set
$\beta(p)=\{x\in\mathbf{B}(X)\mid x(p)=0\}$
is the kernel of a surjective ring homomorphism $\mathbf{B}(X)\to\mathbb{F}_{2}$, so it is a maximal ideal in $\mathbf{B}(X)$. Consequently, we have an induced map
$\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X)).$
For every $x\in\mathbf{B}(X)$ the pre-image
$\beta^{-1}(D(x))=\{p\in X\mid x(p)=1\}$
is open. Since the sets $\{D(x)\mid x\in\mathbf{B}(X)\}$ form a sub-basis of $\mathrm{Spec}(\mathbf{B}(X))$, we get that $\beta\colon X\to\mathrm{Spec}(\mathbf{B}(X))$ is continuous.
Theorem A.8.
When $X$ is compact and totally separated, then $\beta$ is a homeomorphism.
Proof.
Since $X$ is compact and $\mathrm{Spec}(\mathbf{B}(X))$ is Hausdorff, it will be sufficient to show that $\beta$ is a bijection. Let $x,y\in X$ be two distinct points. Since $X$ is totally separated, so there exist disjoint open sets $U\subset X$ containing $x$ and $V\subset X$ containing $y$ such that $X$ is the union of $U$ and $V$. Let $f\colon X\to\mathbb{F}_{2}$ be the characteristic function of $U$. It is in $\mathcal{B}(X)$ since the complement of $U$ is open, too. Clearly $f\in\beta(y)$, but $f\not\in\beta(x)$, so $\beta$ is injective.
For every ideal $I\triangleleft\mathbf{B}(X)$ let $Z(I)\subseteq X$ denote the closed subset
$Z(I)=\{x\in X\mid f(x)=0\quad(\forall f\in I)\}.$
We claim that for every proper ideal $I\triangleleft\mathbf{B}(X)$ the set $Z(I)$ is non-empty. First consider the case when $I=(f)$ for some $f\in\mathbf{B}(X)$. Then
$Z(I)=\{x\in X\mid f(x)=0\},$
so if this set is empty, then $f$ is the identically one function, and hence $I=(1)=\mathcal{B}(X)$ is not proper, a contradiction. Next consider the case when $I$ is finitely generated; then it is principal by part $(iv)$ of Proposition A.3, so $Z(I)$ is non-empty by the above. Finally, consider the general case. Then $Z(I)$ is the intersection of sets of the form $Z(J)$ where $J$ is a finitely generated ideal of $I$. Since the latter collection of sets is closed under finite intersections, and each member is non-empty by the above, we get $Z(I)$ is also non-empty, since $X$ is compact.
Now let $\mathfrak{m}\triangleleft\mathbf{B}(X)$ be a maximal ideal. By the above there is an $x\in Z(\mathfrak{m})$. Clearly $\beta(x)\supseteq\mathfrak{m}$, but $\mathfrak{m}$ is maximal, and hence $\beta(x)=\mathfrak{m}$. Therefore $\beta$ is surjective, too. ∎
Notation A.9.
Let $\mathbf{BO}$ denote the category of Boolean rings where morphisms are ring homomorphisms, and let $\mathbf{CTS}$ denote the category of compact, totally separated topological spaces where morphisms are continuous maps. There are two contravariant functors
$\mathbf{B}\colon\mathbf{CTS}\to\mathbf{BO},\quad X\mapsto\mathbf{B}(X),$
which is well-defined as we saw in Examples A.2, and
$\mathbf{S}\colon\mathbf{BO}\to\mathbf{CTS},\quad R\mapsto\mathrm{Spec}(R),$
well-defined by Proposition A.4.
Theorem A.10 (Stone duality).
The functors $\mathbf{B}$ and $\mathbf{S}$ are a pair of dualities of categories.
Proof.
By Theorem A.6, the map $\sigma$ is a natural isomorphism between the identity of $\mathbf{BO}$ and $\mathbf{B}\circ\mathbf{S}$. By Theorem A.8, the map $\beta$ is a natural isomorphism between the identity of $\mathbf{CTS}$ and $\mathbf{S}\circ\mathbf{B}$. ∎
Corollary A.11.
Let $R$ be a finite Boolean ring. Then the following are equivalent:
  1. $(i)$
    $R$ is finite.
  2. $(ii)$
    $R$ is finitely generated as an $\mathbb{F}_{2}$-algebra.
  3. $(iii)$
    $R$ is Noetherian.
  4. $(iv)$
    $R$ is Artinian.
  5. $(v)$
    $\mathrm{Spec}(R)$ is finite.
  6. $(vi)$
    $R\cong\mathbb{F}_{2}^{X}$ for a finite set $X$.
In this case $R\cong\mathbb{F}_{2}^{\mathrm{Spec}(R)}$.
Proof.
Every Boolean ring is an $\mathbb{F}_{2}$-algebra, so if $R$ is finite, then it is finitely generated as an $\mathbb{F}_{2}$-algebra, and hence $(i)$ implies $(ii)$. Every finitely generated $\mathbb{F}_{2}$-algebra is Noetherian by Hilbert’s basis theorem, so $(ii)$ implies $(iii)$. Every Boolean ring is zero dimensional by part $(ii)$ of Proposition A.3, so if it is Noetherian, it is Artinian by a standard theorem in commutative algebra (see Theorem 8.5 of [1] on page 90). Therefore $(iii)$ implies $(iv)$. Every Artinian ring is a finite direct product of Artinian local rings (see Theorem 8.7 of [1] on page 90), so its spectrum is finite. Therefore $(iv)$ implies $(v)$. Now let $R$ be a Boolean ring whose spectrum is finite. Since $\mathrm{Spec}(R)$ is totally separated by Proposition A.4, it is discrete, and hence $R\cong\mathbb{F}_{2}^{\mathrm{Spec}(R)}$ by Theorem A.6. In particular, $(v)$ implies $(vi)$. If $R\cong\mathbb{F}_{2}^{X}$ for a finite set $X$, then $R$ is clearly finite, so $(vi)$ implies $(i)$. ∎
Notation A.12.
Let $\mathbf{FBO}$ denote the category of finite Boolean rings where morphisms are ring homomorphisms, and let $\mathbf{FTS}$ denote the category of finite, totally separated topological spaces where morphisms are continuous maps. Note that the latter is the same as the category of finite, discrete topological spaces. There are two restrictions of functors
$\mathbf{B}|_{\mathbf{FTS}}\colon\mathbf{FTS}\to\mathbf{FBO},\quad X\mapsto\mathbf{B}(X)$
and
$\mathbf{S}|_{\mathbf{FBO}}\colon\mathbf{FBO}\to\mathbf{FTS},\quad R\mapsto\mathrm{Spec}(R),$
where the latter is well-defined by Corollary A.4.
Theorem A.13.
The functors $\mathbf{B}|_{\mathbf{FTS}}$ and $\mathbf{S}|_{\mathbf{FTS}}$ are a pair of dualities of categories.
Proof.
The restrictions of the natural isomorphisms $\beta$ and $\sigma$ onto $\mathbf{FTS}$ and $\mathbf{FBO}$ are the respective required natural isomorphisms. ∎
Definition A.14.
Let $R$ be a Boolean ring. We say that two elements $x,y\in R$ are orthogonal if $xy=0$. We say that an element $a\in R$ is an atom if it is non-zero and cannot be written as the sum of two non-zero orthogonal elements of $R$. The support of an element $x\in R$ is the subset $D(f)\subseteq\mathrm{Spec}(X)$ introduced in (19).
Lemma A.15.
Let $R$ be a Boolean ring. Then the following holds:
  1. $(i)$
    Two elements $x,y\in R$ are orthogonal if and only if the intersection of their support is empty.
  2. $(ii)$
    A non-zero element $a\in R$ is an atom if and only if its support cannot be written as the disjoint union of two non-empty open and closed subsets.
  3. $(iii)$
    Every pair of different atoms of $R$ are orthogonal to each other.
Proof.
Note that the support of the product $xy$ is the intersection of the supports of $x$ and $y$. Since the support of an element of $R$ is empty if and only if it is zero by Theorem A.6, claim $(i)$ follows. If $a=x+y$ such that $x,y$ are orthogonal and both non-zero, then the support of $a$ is the disjoint union of the supports of $x$ and $y$ by part $(i)$. On the other hand if the support of $a$ is the disjoint union of the non-empty open and closed subsets $X$ and $Y$, there are elements $x,y\in R$ whose support is $X$, $Y$, respectively, by Theorem A.6. By part $(i)$ these are orthogonal, non-zero and their sum has the same support as $a$. So $a=x+z$, and hence claim $(ii)$ is true.
Let $a,b\in R$ be two atoms whose product is non-zero. Then $a=ab+a(1-b)$ and $aba(1-b)=a^{2}(b-b^{2})=0$, so $a(1-b)=0$ by the definition of atoms. We get that $a=ab$. The same reasoning for $b$ show that $b=ab$. Therefore $a=b$, and hence $(iii)$ holds. ∎
Corollary A.16.
Let $R$ be a finite Boolean ring. Then the atoms of $R$ form a natural basis of $R$ whose elements are orthogonal to each other.
Proof.
By Corollary A.11 the topological space $\mathrm{Spec}(R)$ is finite, discrete, and $R\cong\mathbb{F}_{2}^{\mathrm{Spec}(R)}$. Therefore an element of $R$ is an atom if and only if its support is a point by claim $(ii)$ of Lemma A.15. These clearly form a basis of $R$ and they are orthogonal by claim $(iii)$ of Lemma A.15. ∎
Appendix B Proof of Proposition 5.9
For lack of a reference known to the authors, we provide a proof of the existence of the spectral sequence claimed in Proposition 5.9. The proof is an adjustment of the construction of a spectral sequence for the composition of a right derived functor with an inductive limit in [15, Chapter 4]. Even though Hochschild cohomology is a right derived functor we need to adjust for the fact that we take Hochschild cohomology of a system over varying base rings.
Proof of Proposition 5.9.
We recall from Remark 4.4 and (12) that we can identify the graded Hochschild cohomology of $A$ and $M$ with the cohomology
$\mathrm{HH}^{n,s}(A,M)=H^{n}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes*},M[s]),d^{*}).$
To compute the higher inverse limits occurring in the spectral sequence we will use the resolutions of [15, Chapter 4]. Let $\{S_{\alpha},f_{\alpha\beta}\}_{\alpha}$ be a projective system of $\mathbb{F}_{2}$-vector spaces. We denote by $(\Pi^{*}S_{\alpha},\delta^{*})$ the cochain complex defined in [15, Chapter 4, page 32] whose $n$th space $\Pi^{n}S_{\alpha}$ is the product
$\prod_{\alpha_{0}\leq\ldots\leq\alpha_{n}}S_{\alpha_{0}\ldots\alpha_{p}}~{}\text{with}~{}S_{\alpha_{0}\ldots\alpha_{n}}=S_{\alpha_{0}}$
consisting of families $s_{\alpha_{0}\ldots\alpha_{n}}$ of elements in $S_{\alpha_{0}}$ indexed over tuples $\alpha_{0}\leq\ldots\leq\alpha_{n}$ in $I$. These families inherit the structure of $\mathbb{F}_{2}$-vector spaces with component-wise operations. By [15, Théorème 4.1], the $p$th higher inverse limit of the system $\{S_{\alpha},f_{\alpha\beta}\}$ equals the $p$th cohomology group of this cochain complex, i.e.,
[…]
For an inductive system $\{T_{\alpha},g_{\alpha\beta}\}_{\alpha}$, we denote by $(\Sigma^{*}T_{\alpha},\partial^{*})$ the chain complex defined in [15, page 33] with
$\Sigma^{n}T_{\alpha}=\bigoplus_{\alpha_{0}\leq\ldots\leq\alpha_{n}}T_{\alpha_{0}\ldots\alpha_{n}}~{}\text{with}~{}T_{\alpha_{0}\ldots\alpha_{n}}=T_{\alpha_{0}}.$
The chain complex $\Sigma^{*}T_{\alpha}$ is dual to $\Pi^{*}$ in the sense that
(20) $\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(\Sigma^{*}T_{\alpha},M)\cong\Pi^{*}\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(T_{\alpha},M).$
Now we construct the spectral sequence following [15, page 34–35]. We denote the transition maps of the directed system of subrings $A_{\alpha}$ of $A$ by $\iota_{\alpha\beta}\colon A_{\beta}\hookrightarrow A_{\alpha}$. We define a double complex $C^{*,*}=\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(\Sigma A_{\alpha},M)$ by setting $C^{p,q}$ to be
$C^{p,q}=\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}\left(\bigoplus_{\alpha_{0}\leq\ldots\leq\alpha_{p}}(A_{\alpha_{0}\ldots\alpha_{p}})^{\otimes q},M\right).$
The differential in the $p$-direction is the one induced by $\partial^{*}$. For fixed $p$ and $A_{\alpha_{0}\ldots\alpha_{p}}=A_{\alpha_{0}}$, the differential in the $q$-direction is the differential $d^{*}$ of the complex $\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}((A_{\alpha_{0}\ldots\alpha_{p}})^{\otimes q},M)$.
We consider the two spectral sequences that arise from this double complex. First let $p$ be the filtration-index. Then, using isomorphism (20), we get
${E^{\prime}_{1}}^{p,q}$ $=\prod_{\alpha_{0}\leq\ldots\leq\alpha_{p}}H^{q}\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}((A_{\alpha_{0}\ldots\alpha_{p}})^{\otimes*},M)$
$=\prod_{\alpha_{0}\leq\ldots\leq\alpha_{p}}\mathrm{HH}^{q,s}(A_{\alpha_{0}\ldots\alpha_{p}},M).$
By [15, Théorème 4.1] applied to the projective system $\{\mathrm{HH}^{q,s}(A_{\alpha},M),\mathrm{HH}^{q,s}(f_{\alpha\beta},M)\}_{\alpha}$, taking cohomology in the $p$-direction gives
[…]
Now let $q$ be the filtration-index. Taking cohomology in the $p$-direction gives
${E^{\prime\prime}_{1}}^{p,q}=H^{p}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(\Sigma^{*}A_{\alpha},M[s]).$
[…]
Since inductive limits commute with tensor products, we have
[…]
for every $q$. Finally, taking cohomology in the direction of $q$ gives
${E^{\prime\prime}_{2}}^{p,q}=\begin{cases}H^{q}(\underline{\mathrm{Hom}}_{\mathbb{F}_{2}}(A^{\otimes*},M[s]))=\mathrm{HH}^{q,s}(A,M)&\text{if}~{}p=0\\ 0&\text{if}~{}p>0.\end{cases}$
This spectral sequence degenerates. Hence we get that there is a spectral sequence of the form
[…]